### v1

# Non-Asymptotic Analysis of Efficiency in Conformalized Regression

Abstract
Conformal prediction provides prediction sets with coverage guarantees. The informativeness of conformal prediction depends on its efficiency, typically quantified by the expected size of the prediction set. Prior work on the efficiency of conformalized regression commonly treats the miscoverage level $\alpha$ as a fixed constant. In this work, we establish non-asymptotic bounds on the deviation of the prediction set length from the oracle interval length for conformalized quantile and median regression trained via SGD, under mild assumptions on the data distribution. Our bounds of order $\mathcal{O}(1/\sqrt{n}+1/(\alpha^{2}n)+1/\sqrt{m}+\exp(-\alpha^{2}m))$ capture the joint dependence of efficiency on the proper training set size $n$, the calibration set size $m$, and the miscoverage level $\alpha$. The results identify phase transitions in convergence rates across different regimes of $\alpha$, offering guidance for allocating data to control excess prediction set length. Empirical results are consistent with our theoretical findings.
Keywords: conformal prediction, efficiency, conformalized regression, quantile regression, uncertainty quantification
1 Introduction
Deploying machine learning models in safety-critical domains, such as health care (Allgaier et al., 2023; Gui et al., 2024), finance (Wisniewski et al., 2020; Bastos, 2024), and autonomous systems (Lindemann et al., 2023; Ren et al., 2023), requires not only accurate predictions but also reliable uncertainty quantification. Conformal prediction (CP) is a principled, distribution-free framework for this purpose, equipping black-box models with prediction sets achieving coverage guarantees or validity (Vovk et al., 2005; Balasubramanian et al., 2014). Formally, given a set of data $\{(X_{j},Y_{j})\}_{j=1}^{m}$ drawn from a distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$, for any user-specified miscoverage level $\alpha\in(0,1)$ and a predictive model, conformal prediction constructs a set-valued function $\mathcal{C}:\mathcal{X}\to 2^{\mathcal{Y}}$ such that, for a test pair $(X_{m+1},Y_{m+1})\sim\mathcal{P}$, the prediction set $\mathcal{C}(X_{m+1})$ covers the label $Y_{m+1}$ with probability
$\mathbb{P}\left[Y_{m+1}\in\mathcal{C}(X_{m+1})\right]\geq 1-\alpha.$ (1)
Split conformal prediction is a computationally efficient variant that incorporates training predictive models. It splits data into a proper training set and a calibration set; the model is first trained on the former, and its uncertainty is then quantified using the latter. During calibration, nonconformity score functions are constructed to measure the discrepancy between model predictions and true labels. The distribution of these scores is estimated over the calibration set, and a quantile of them defines a threshold. The prediction set $\mathcal{C}$ is then obtained by collecting all candidate labels whose nonconformity scores are no larger than this threshold.
A central focus of conformal prediction is efficiency, commonly quantified by the expected measure of the prediction set (Shafer and Vovk, 2008). For classification tasks, efficiency relates to the cardinality of the predicted label set; for regression, it corresponds to the length (or volume) of the prediction interval (or region). Under the validity condition (1), smaller prediction sets are more informative. Early works primarily evaluated efficiency empirically, whereas recent research has shifted toward asymptotic efficiency, demonstrating that prediction sets converge to the oracle sets as the sample size increases (Sesia and Candès, 2020; Chernozhukov et al., 2021; Izbicki et al., 2022). In contrast, non-asymptotic efficiency, or finite-sample guarantees on the expected measure or excess measure of the prediction set, remains much less understood, with only partial results available (Lei and Wasserman, 2014; Lei et al., 2018; Dhillon et al., 2024; Bars and Humbert, 2025). Existing non-asymptotic bounds are typically expressed based on the calibration set size $m$, whereas the effect of training set size $n$ and miscoverage level $\alpha$ remains an open question in split conformalized regression.
In this work, we analyze the efficiency of split conformal prediction in regression, focusing on conformalized median regression (CMR) and conformalized quantile regression (CQR) (Romano et al., 2019). CMR uses the absolute residual as the nonconformity score, and the quantile of the calibration residuals then determines the half-width of a symmetric prediction interval centered at the estimated conditional median. In contrast, CQR estimates both upper and lower conditional quantiles, defining nonconformity scores relative to these estimates. After calibration, CQR yields adaptive, asymmetric prediction intervals that naturally capture heteroscedasticity without assuming symmetric conditional quantiles.
Contributions.
We present a non-asymptotic theoretical analysis of the efficiency of conformalized quantile regression and conformalized median regression under stochastic gradient descent (SGD) training. Our main contributions are as follows:
  • Finite-sample bounds for CQR. For CQR-SGD (Algorithm 1), we derive an upper bound of order $\mathcal{O}(1/\sqrt{n}+1/(\alpha^{2}n)+1/\sqrt{m}+\exp(-\alpha^{2}m))$ on the expected deviation of the prediction set length from the oracle interval, where $n$ is the proper training set size, $m$ is the calibration set size, and $\alpha$ is the miscoverage level (Theorem 3.2). Unlike prior work that relies on assumptions on intermediate quantities, our analysis places assumptions directly on the data distribution.
  • Finite-sample bounds for CMR. For homoscedastic tasks, CMR-SGD produces symmetric intervals of constant length across inputs, enabling us to derive a non-asymptotic upper bound of analogous order (Theorem 4.1) to CQR.
  • Theoretical guidance. To the best of our knowledge, our work is the first analysis establishing upper bounds on interval length deviation as a function of $(n,m,\alpha)$, revealing phase transitions across different $\alpha$ regimes (Section 3.2.1). Our results thus offer guidance on allocating data between training and calibration to control excess length at a desired miscoverage level. These theoretical insights are further validated through experiments.
Finally, while our theorems are presented for models trained with SGD, the analytical framework developed in this paper is not tied to a specific optimizer: the bounds extend directly to other optimization algorithms by substituting their corresponding estimation error rates.
2 Preliminaries
Quantiles of random variables.
For $\gamma\in\left(0,1\right)$, the $\gamma$-quantile of a random variable $Z$ with cumulative distribution function (c.d.f.) $F$ is defined as the set
$\mathcal{Q}_{\gamma}(Z):=\left\{u\in\mathbb{R}:F\left(u\right)\geq\gamma\text{ and }F\left(u^{-}\right)\leq\gamma\right\}$
where $F\left(u^{-}\right)$ denotes the left limit of $F$ at $u$. A canonical representative is
$q_{\gamma}(Z):=\inf\{u\in\mathbb{R}:F(u)\geq\gamma\}.$
In the case where $F$ is continuous and strictly increasing at $q_{\gamma}(Z)$, the quantile set reduces to a singleton, i.e., $\mathcal{Q}_{\gamma}(Z)=\left\{q_{\gamma}(Z)\right\}$.
Conditional quantile function.
For $(X,Y)\sim\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$, the conditional $\gamma$-quantile function $q_{\gamma}\left(Y\mid X\right):\mathcal{X}\to\mathbb{R}$ is defined as
$q_{\gamma}\left(Y\mid X=x\right):=\inf\left\{u\in\mathbb{R}:F_{Y\mid X=x}\left(u\right)\geq\gamma\right\}\quad\text{for all }x\in\mathcal{X}$ (2)
Split conformal prediction.
In split conformal prediction, the data are partitioned into the proper training set $\mathcal{D}_{\text{train}}$ and the calibration set $\mathcal{D}_{\text{cal}}$. The training set is first used to train a model $h$. With the trained model $h$, a nonconformity score function $\psi_{h}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ is then defined to quantify the discrepancy between a candidate label $y$ and the input $x$, where higher scores indicate worse conformity. The nonconformity scores $S_{m}:=\{\psi_{h}(x_{j},y_{j})\}_{j=1}^{m}$ are computed for all calibration samples in $\mathcal{D}_{\mathrm{cal}}=\{(x_{j},y_{j})\}_{j=1}^{m}$. The sample quantile $\hat{q}_{(1-\alpha)_{m}}$ is calculated at level:
$(1-\alpha)_{m}\;:=\;{\lceil(1-\alpha)(m+1)\rceil}\ /\ {m},$
corresponding to the $\lceil(1-\alpha)(m+1)\rceil$-th smallest value in $S_{m}$, which is also known as the empirical quantile. The prediction set for a new input $x$ is then defined as
$\mathcal{C}(x)\;=\;\{\,y\in\mathcal{Y}:\psi_{h}(x,y)\leq\hat{q}_{(1-\alpha)_{m}}\,\}.$
Bachmann–Landau notation.
We employ Bachmann–Landau (or Big O) notation in the limit as $n,m\to\infty$. For positive sequences or functions $f,g$, we write $f=O(g)$ if there exists $C,N>0$ such that $|f(k)|\leq C\ |g(k)|$ for all $k\geq N$; we write $f=\Omega(g)$ if there exists c,N>0\,c,N>0 such that $|f(k)|\geq c\,|g(k)|$ for all $k\geq N$. We write $f=o(g)$ if $f/g\to 0$, and $f=\omega(g)$ if $f/g\to\infty$.
3 Analysis of Conformalized Quantile Regression (CQR)
3.1 Problem Setup for CQR-SGD
Data model.
We consider a random design setting where training, calibration, and test samples are drawn i.i.d. from an unknown distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$. Formally, for all $i\in[n],j\in[m]$
$(X_{i}^{\mathrm{train}},Y_{i}^{\mathrm{train}}),\;(X_{j}^{\mathrm{cal}},Y_{j}^{\mathrm{cal}}),\;(X^{\mathrm{test}},Y^{\mathrm{test}})\quad\text{i.i.d.}\;{\sim}\;\;\mathcal{P}.$
We assume the covariate space $\mathcal{X}\subset\mathbb{R}^{d}$ is bounded: there exists a constant $B>0$ such that
$\|x\|_{2}\leq B,\quad\forall\,x\in\mathcal{X}.$ (3)
Similarly, the response space $\mathcal{Y}\subset\mathbb{R}$ is assumed to be a bounded interval $y_{\min},y_{\max}]$.
Learning objective.
In CQR, the training set $\mathcal{D}_{\mathrm{train}}$ is used to estimate the conditional $\gamma$-quantile function $q_{\gamma}\left(Y\mid X\right)$ defined in (2), where $\gamma=1-\alpha/2,\alpha/2$. The estimated function $t_{\gamma}(\cdot;\theta_{n}(\gamma))$ is obtained by solving the stochastic pinball loss minimization problem (Koenker and Bassett Jr, 1978):
$\min_{\theta\in\Theta}\;\ell_{\gamma}(\theta):=\mathbb{E}_{(X,Y)\sim\mathcal{P}_{X\times Y}}\bigl[\,L_{\gamma}\bigl(t_{\gamma}(X;\theta),\,Y\bigr)\bigr],$ (4)
where the pinball loss takes the form
$L_{\gamma}(t,y)=\gamma(y-t)\,\mathbf{1}\{y\geq t\}+(1-\gamma)(t-y)\,\mathbf{1}\{y<t\}.$ (5)
We consider a linear function class with a convex and compact parameter space:
$t_{\gamma}(x;\theta)=\theta^{\top}x,\quad\theta\in\Theta\subset\mathbb{R}^{d},\quad\sup_{\theta\in\Theta}\|\theta\|_{2}\leq K<\infty,$ (6)
Without loss of generality, we assume $K\leq\max\{|y_{\min}|,|y_{\max}|\}/B$. The linear model represents a standard setting for theoretical analysis of quantile regression (Koenker, 2005; Pan and Zhou, 2021), ensuring convexity of the objective function in (4).
Learning algorithm.
To solve (4), we consider the stochastic approximation framework (Robbins and Monro, 1951), focusing on stochastic gradient descent (SGD). The $\theta$ is updated according to
$\theta_{k+1}=\Pi_{\Theta}\!\left(\theta_{k}-\eta_{k}\hat{g}_{k}\right),$ (7)
where $\eta_{k}$ is the step size, $\Pi_{\Theta}$ denotes the Euclidean projection onto $\Theta$, and $\hat{g}_{k}$ is a stochastic subgradient satisfying $\mathbb{E}[\hat{g}_{k}\mid\theta_{k}]=g_{k}$, with $g_{k}$ a subgradient of the population objective in(4) at $\theta_{k}$.
Let $\theta_{n}(\gamma)$ denote the parameter learned by solving (4) via SGD on the training set $\mathcal{D}_{\mathrm{train}}$. For convenience, we introduce the shorthand notations for the learned parameters
$\underline{\theta}_{n}:=\theta_{n}\!\left(\alpha/2\right),\quad\bar{\theta}_{n}:=\theta_{n}\!\left(1-\alpha/2\right),\quad\vartheta_{n}:=\left(\underline{\theta}_{n},\bar{\theta}_{n}\right).$
Conformalized quantile regression.
CQR employs two estimated conditional quantile functions, $t_{\alpha/2}(\cdot;\underline{\theta}_{n})$ and $t_{1-\alpha/2}(\cdot;\bar{\theta}_{n})$. Given the learned parameters $\vartheta_{n}=\left(\underline{\theta}_{n},\bar{\theta}_{n}\right)$, the score for $(X,Y)$ is
$S\left(X,Y;\vartheta_{n}\right):=\max\bigl\{\,t_{\alpha/2}(X;\underline{\theta}_{n})-Y,\;\;Y-t_{1-\alpha/2}(X;\bar{\theta}_{n})\,\bigr\}.$ (8)
Thus $S>0$ if $Y$ lies outside the interval $t_{\alpha/2}(X;\underline{\theta}_{n})),t_{1-\alpha/2}(X;\bar{\theta}_{n})]$, and $S\leq 0$ otherwise. Let $S_{m}(\mathcal{D}_{\mathrm{cal}};\vartheta_{n})$ denote the $m$ scores on the calibration data, and let $\hat{q}_{(1-\alpha)_{m}}(S_{m}\mid\vartheta_{n})$ be their empirical $(1-\alpha)_{m}$-quantile, i.e., the $\lceil(1-\alpha)(m+1)\rceil$-th smallest value of $S_{m}(\mathcal{D}_{\mathrm{cal}};\vartheta_{n})$. The prediction set for a test covariate $X$ is then
$\mathcal{C}(X)=\left[t_{\alpha/2}\left(X;\underline{\theta}_{n}\right)-\hat{q}_{(1-\alpha)_{m}}(S_{m}\mid\vartheta_{n}),\;t_{1-\alpha/2}\left(X;\bar{\theta}_{n}\right)+\hat{q}_{(1-\alpha)_{m}}(S_{m}\mid\vartheta_{n})\right],$ (9)
if $t_{1-\alpha/2}\left(X;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(X;\bar{\theta}_{n}\right)+2\hat{q}_{(1-\alpha)_{m}}(S_{m}\mid\vartheta_{n})\geq 0$; otherwise, $\mathcal{C}(X)=\emptyset$.
Remark 3.1.
The phenomenon where the lower quantile estimate exceeds the upper quantile estimate is known as quantile crossing (Romano et al., 2019; Bassett Jr and Koenker, 1982). We show in the proof of Proposition B.8 that, quantile crossing does not occur with high probability once the training set size $n$ is sufficiently large.
The whole pipeline of CQR with SGD training is summarized in Algorithm 1.
Algorithm 1 Conformalized Quantile Regression with SGD Training (CQR-SGD)
1:Input: Dataset of size $(n+m)$, miscoverage level $\alpha$, new input $x$
2: Split the dataset into a proper training set $\mathcal{D}_{\mathrm{train}}$ of size $n$ and a calibration set $\mathcal{D}_{\mathrm{cal}}$ of size $m$
3: Train quantile regressors $t_{\alpha/2}(\cdot;\underline{\theta}_{n})$ and $t_{1-\alpha/2}(\cdot;\bar{\theta}_{n})$ on $\mathcal{D}_{\mathrm{train}}$ by solving (4) via SGD
4: Compute $m$ nonconformity scores on $\mathcal{D}_{\mathrm{cal}}$ according to (8)
[…]
7:Output: Prediction set $\mathcal{C}(x)$ for a new input $x$
3.2 Theoretical Results for Efficiency of CQR
To establish upper bounds on the expected length deviation of the prediction sets, we introduce the following assumptions.
Assumption 3.1 (Well-specification in CQR).
For $\gamma\in\{\alpha/2,1-\alpha/2\}$, there exists $\theta^{*}(\gamma)\in\Theta$ such that
$q_{\gamma}(Y\mid X=x)=t_{\gamma}(x;\theta^{*}(\gamma))\quad\text{for all }x\in\mathcal{X}.$
Assumption 3.1 ensures that $\theta^{*}(\gamma)$ is a minimizer of (4) (Takeuchi et al., 2006; Steinwart and Christmann, 2011).
Similar to $\underline{\theta}_{n},\bar{\theta}_{n}$, and $\vartheta_{n}$, we introduce the shorthand notations for the ground-truth parameters
$\underline{\theta}^{*}:=\theta^{*}\!\left(\alpha/2\right),\quad\bar{\theta}^{*}:=\theta^{*}\!\left(1-\alpha/2\right),\quad\vartheta^{*}:=\left(\underline{\theta}^{*},\bar{\theta}^{*}\right).$
Assumption 3.2 (Bounded covariance).
There exist constants $0<\lambda_{\min}\leq\lambda_{\max}<\infty$ such that
$\lambda_{\min}I\;\preceq\;\Sigma:=\mathbb{E}[XX^{\top}]\;\preceq\;\lambda_{\max}I,$ (10)
where $I$ is the identity matrix, and $A\preceq B$ means that $\left(B-A\right)$ is positive semi-definite for two symmetric matrices $A,B$.
Note that $\lambda_{\max}\leq B^{2}$, since $\|x\|_{2}\leq B$ for all $x\in\mathcal{X}$.
Assumption 3.3 (Regularity of the conditional density).
For any $x\in\mathcal{X}$, the conditional probability density function (p.d.f.) $f_{Y\mid X}(\,\cdot\mid x)$ exists and is continuous. Moreover, there exist constants $0<f_{\min}\leq f_{\max}<\infty$ such that
$f_{\min}\;\leq\;f_{Y\mid X}(y\mid x)\;\leq\;f_{\max},\qquad\forall\,x\in\mathcal{X},\;\;\forall\,y\in\mathcal{Y}.$ (11)
We notice that Assumption 3.3 concerns only the underlying data distribution $\mathcal{P}$. In particular, our assumptions are agnostic to the induced nonconformity scores, unlike prior works which impose assumptions on the induced distribution of nonconformity scores, which depends on the trained predictive model. Assumption 3.3 is satisfied by many common continuous distributions once truncated to a bounded support and normalized, including the truncated normal distribution.
Assumption 3.3 implies that the conditional support of $Y$ given any $x\in\mathcal{X}$ is the common set $\mathcal{Y}$. The lower bound $f_{Y\mid X}(y\mid x)\geq f_{\min}$ guarantees that $\mathcal{Y}$ is bounded, while the upper bound $f_{Y\mid X}(y\mid x)\leq f_{\max}$ ensures that $\mathcal{Y}$ has non-empty interior. A constant $H$ is defined to characterize the flatness of conditional distribution, i.e.
$H(f_{\max},f_{\min}):=f_{\max}\ /\ f_{\min}.$ (12)
In particular, the Lebesgue measure of $\mathcal{Y}$ satisfies $1/f_{\max}\leq|\mathcal{Y}|\leq 1/f_{\min}$. Together with $B$ in (3), $K$ in (6), and Assumption 3.1, it yields
$|y|\;\leq\;BK+{1}/{f_{\min}},\qquad\forall\,y\in\mathcal{Y}.$ (13)
The score $S$ has a bounded support, since $|t_{1/2}(X;\theta_{n})|\leq BK$ and $|Y|\leq BK+1/f_{\min}$, i.e.,
$|S|\leq R:=2BK+{1}/{f_{\min}}.$
As a first step toward bounding the expected length deviation, Theorem 3.1 establishes upper bounds on both the prediction error of the quantile regressor and the parameter estimation error under SGD training, expressed in terms of the training sample size $n$.
Theorem 3.1 (Quantile regression error of SGD-trained models).
If Assumptions 3.13.3 hold, then
$\mathbb{E}_{X,\theta_{n}}\left[\left(t_{\gamma}\left(X;\theta_{n}\left(\gamma\right)\right)-t_{\gamma}\left(X;\theta^{*}\left(\gamma\right)\right)\right)^{2}\right]$ $\leq\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{3}f_{\min}^{2}n},$ (14)
$\mathbb{E}_{\theta_{n}}\left[\|\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\|_{2}^{2}\right]$ $\leq\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{4}f_{\min}^{2}n}.$ (15)
The proof of Theorem 3.1 is deferred to Appendix B.1.
Remark 3.2.
The results of Theorem 3.1 are established under a strongly-convex assumption as they rely on Theorem B.4 from Rakhlin et al. (2012). Comparable rates can also be obtained for non-strongly-convex objectives under the assumptions in Bach and Moulines (2013), where Assumption 3.2 can be weakened to requiring only the invertibility of $\mathbb{E}[XX^{\top}]$.
Remark 3.3.
Faster rates than those of Theorem 3.1 are attainable with variance-reduced stochastic gradients; see Appendix A for further discussion.
Theorem 3.2 establishes a non-asymptotic efficiency guarantee for CQR-SGD (Algorithm 1), bounding the expected length deviation of the prediction set from the oracle conditional quantile interval
$\mathcal{C}^{*}(X):=\left[\ q_{\alpha/2}\left(Y\mid X\right),\;q_{1-\alpha/2}\left(Y\mid X\right)\ \right].$ (16)
We measure the efficiency of conformalized regression methods by the expected length deviation
$\mathbb{E}_{X,\vartheta_{n},\mathcal{D}_{\mathrm{cal}}}\bigl[\bigl|\left|\mathcal{C}(X)\right|-\left|\mathcal{C}^{*}(X)\right|\bigl|\bigl].$ (expected length deviation)
Theorem 3.2 (Efficiency of CQR-SGD).
For CQR-SGD, suppose Assumptions 3.13.3 hold. If $m>{8H}/{\min\{\alpha,1-\alpha\}}$, then for test sample $(X,Y)$ and $0<\alpha\leq 1/2$, $\mathbb{E}_{X,\vartheta_{n},\mathcal{D}_{\mathrm{cal}}}\bigl[\bigl|\left|\mathcal{C}(X)\right|-\left|\mathcal{C}^{*}(X)\right|\bigl|\bigl]\leq\mathcal{O}\!\left(n^{-1/2}+(\alpha^{2}n)^{-1}+m^{-1/2}+\exp(-\alpha^{2}m)\right)$ (17) where $H$ is the constant defined in (12).
The explicit upper bound (B.5) and the full proof of Theorem 3.2 are presented in Appendix C, with a proof sketch illustrated in Figure 1.
[…]
Figure 1: Proof outline of Theorem 3.2. Full proof deferred to Section B.
Remark 3.4.
While Theorem 3.2 is presented for CQR trained using SGD, the analysis strategy applies to other optimization algorithms. In particular, one can replace the SGD error bound in Theorem 3.1 with that of the chosen optimizer. This replacement modifies only the terms in the overall bound that depend on the training set size $n$. Formally, suppose the upper bound in Theorem 3.1 is replaced by $\varphi_{n}$ where $\varphi_{n}\to 0$ as $n\to\infty$, then the upper bound in Theorem 3.2 becomes $\mathcal{O}\left(\varphi_{n}^{1/2}+\alpha^{-2}\varphi_{n}+m^{-1/2}+\exp(-\alpha^{2}m)\right)$.
Remark 3.5.
For a random variable $Z$, the density level set $\mathcal{L}(u_{1-\alpha})$ is the optimal prediction set with coverage probability $1-\alpha$ (Lei et al., 2011), i.e.,
$\mathcal{L}(u_{1-\alpha}):=\left\{z\in\mathcal{Z}:f_{Z}(z)\geq u_{1-\alpha}\right\}=\operatorname*{arg\,min}_{\mathbb{P}[Z\in\mathcal{C}]\geq 1-\alpha}|\mathcal{C}|$
where $u_{1-\alpha}=\inf\{u:\mathbb{P}[Z\in\mathcal{L}(u)]\geq 1-\alpha\}$. The oracle interval $\mathcal{C}^{*}(x)$ coincides with the optimal prediction set if for any $y\in\mathcal{C}^{*}(x)$ and any $y^{\prime}\in\mathcal{Y}\setminus\mathcal{C}^{*}(x)$, it holds that $f_{Y|X=x}(y)\geq f_{Y|X=x}(y^{\prime})$.
3.2.1 Phase Transitions of the Upper Bound
In Theorem 3.2, the upper bound on the expected absolute deviation between the prediction set length $|\mathcal{C}(X)|$ and the oracle interval length $|\mathcal{C}^{*}(X)|$ is expressed explicitly as a function of the training size $n$, calibration size $m$, and miscoverage level $\alpha$. Unlike prior analyses that treat $\alpha$ as a fixed constant, our result reveals its critical role in efficiency. Specifically, the terms $(\alpha^{2}n)^{-1}$ and $\exp(-\alpha^{2}m)$ in the bound imply a fundamental scaling relationship as follows.
Regimes of $\alpha$ in general cases.
  • The length deviation converges to zero whenever $\alpha$ decays slower than $n^{-1/2}$ and $m^{-1/2}$, i.e., $\alpha=\omega(\max\{n^{-1/2},m^{-1/2}\})$. Thus, Theorem 3.2 implies that if the expected prediction set length is required to remain within a fixed tolerance of the oracle length, $\alpha$ is not supposed to be chosen arbitrarily small.
  • For the two $n$-dependent terms in (17), if $\alpha=\Omega(n^{-1/4})$, then they are of order $\mathcal{O}(n^{-1/2})$; otherwise they are of order $\mathcal{O}\left((\alpha^{2}n)^{-1}\right)$.
  • For the two $m$-dependent terms, if $\alpha=\Omega(\sqrt{\log m/m})$, then they are of order $\mathcal{O}(m^{-1/2})$; otherwise they are of order $\mathcal{O}(\exp(-\alpha^{2}m))$.
  • Thus, if $\alpha=$ $\Omega(\max\{n^{-1/4},\sqrt{\log m/m}\})$, the upper bound scales as $\mathcal{O}(n^{-1/2}+m^{-1/2})$, which coincides with the rate in Bars and Humbert (2025) assuming a finite function class.
Regimes of $\alpha$ when $n,m$ of the same order. When $n=\Theta(m)$, the upper bound simplifies to $\mathcal{O}(n^{-1/2}+(\alpha^{2}n)^{-1})$. Figure 2 shows it in different regimes of $\alpha=\Omega(n^{-1})$, consistent with the assumption $m>8H/\min\{\alpha,1-\alpha\}$ in Theorem 3.2.
Data Allocation. If $\alpha=$ $\Omega(\max\{n^{-1/4},\sqrt{\log m/m}\})$, the bound reduces to $\mathcal{O}(n^{-1/2}+m^{-1/2})$, so a natural choice is to set $n$ and $m$ to be of the same order. If $\alpha=$ $\Omega(\sqrt{\log m/m})$ and $\alpha=\omega(n^{-1/4})$, the trade-off is between $\mathcal{O}(m^{-1/2})$ and $\mathcal{O}(1/(\alpha n^{2}))$, and balancing them yields $m=\Theta(\alpha^{4}n^{4})$.
Figure 2: Upper bound orders in Theorem 3.2 in different regimes of $\alpha$ when $n=\Theta(m)$. Results in Lei et al. (2018); Bars and Humbert (2025) lie in the right most regime (blue).
4 Analysis of Conformalized Median Regression (CMR)
4.1 Problem Setup for CMR-SGD
For conformalized median regression (CMR), we consider the same i.i.d. data model and learning algorithm (SGD) as CQR in Section 3.1.
Learning objective.
In CMR, the training set $\mathcal{D}_{\mathrm{train}}$ is used to estimate the conditional median function $q_{1/2}\left(Y\mid X\right)$, which is the special case for conditional $\gamma$-quantile estimation with $\gamma=1/2$ (see (2)). The estimated conditional median function $t_{1/2}(\cdot;\theta)$ is learned by solving the minimization of the expected absolute error (stochastic pinball loss with $\gamma=1/2$) via SGD:
$\min_{\theta\in\Theta}\;\ell_{1/2}(\theta):=\mathbb{E}_{(X,Y)\sim\mathcal{P}_{X\times Y}}\bigl[\,\lvert t_{1/2}(X;\theta)-Y\rvert\,\bigr$]. (18)
We adopt the same linear model class as in CQR, namely (6).
The shorthand notations for the learned parameter $\theta_{n}\!\left(1/2\right)$ and the true parameter $\theta^{*}\!\left(1/2\right)$ are:
$\check{\theta}_{n}:=\theta_{n}\!\left(1/2\right),\quad\check{\theta}^{*}:=\theta^{*}\!\left(1/2\right).$
Conformalized median regression.
In CMR, given the trained regressor $t_{1/2}(\cdot;\check{\theta}_{n})$, the nonconformity score for $(X,Y)$ is
$S\left(X,Y;\check{\theta}_{n}\right):=\left|t_{1/2}(X;\check{\theta}_{n})-Y\right|$ (19)
which corresponds to the absolute prediction error of the estimated conditional median $t{1/2}(\cdot;\check{\theta}_{n})$.
For the calibration set $\mathcal{D}_{\mathrm{cal}}$, let $S_{m}(\mathcal{D}_{\mathrm{cal}};\check{\theta}_{n})$ denote the $m$ scores on calibration data, and let $\hat{q}_{(1-\alpha)_{m}}(S_{m}\mid\check{\theta}_{n})$ be the empirical $(1-\alpha)_{m}$-quantile of $S$ given $\check{\theta}_{n}$, i.e., the $\lceil(1-\alpha)(m+1)\rceil$-th smallest element in $S_{m}(\mathcal{D}_{\mathrm{cal}};\check{\theta}_{n})$. The prediction set for a test covariate $X$ is then
$\mathcal{C}\left(X\right)=\left[t_{1/2}(X;\check{\theta}_{n})-\hat{q}_{(1-\alpha)_{m}}(S_{m}\mid\check{\theta}_{n}),\;\ t_{1/2}(X;\check{\theta}_{n})+\hat{q}_{(1-\alpha)_{m}}(S_{m}\mid\check{\theta}_{n})\right].$ (20)
4.2 Theoretical Results for Efficiency of CMR
The well-specification assumption in CMR assumes a linear $q_{1/2}$:
Assumption 4.1 (Well-specification in CMR).
There exists $\theta^{*}(1/2)\in\Theta$ such that
$q_{1/2}(Y\mid X=x)=t_{1/2}(x;\theta^{*}(1/2))\quad\text{for all }x\in\mathcal{X}.$
For the CMR setting, we make an additional assumption on top of Assumptions 4.1, 3.2, and 3.3:
Assumption 4.2 (Symmetry of quantiles).
There exists $\zeta>0$ such that for every $x\in\mathcal{X}$,
$q_{1-\alpha/2}(Y\mid X=x)-q_{1/2}(Y\mid X=x)=q_{1/2}(Y\mid X=x)-q_{\alpha/2}(Y\mid X=x)=\zeta.$ (21)
Remark 4.1.
Assumption 4.2 is standard in the analysis of conformalized regression based on a single regressor, following the precedent set by Assumption A1 of Lei et al. (2018).
Theorem 4.1 (Efficiency of CMR).
For CMR-SGD, suppose Assumption 4.1,3.2,3.3,4.2 hold. If $m>{8H}/{\min\{\alpha,1-\alpha\}}$, then for test sample $(X,Y)$ and $0<\alpha\leq 1/2$,
$\mathbb{E}_{X,\vartheta_{n},\mathcal{D}_{\mathrm{cal}}}\bigl[\bigl||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)|\bigl|\bigl]\leq\mathcal{O}\!\left(n^{-1/2}+(\alpha^{2}n)^{-1}+m^{-1/2}+\exp(-\alpha^{2}m)\right)$ (22)
where $H$ is the constant defined in (12).
The explicit upper bound (C) and the full proof of Theorem 4.1 are presented in Appendix C.
5 Related Works
Quantile regression.
Quantile regression has attracted significant attention since the seminal work of Koenker and Bassett Jr (1978) due to its robustness to outliers and ability to capture distributional heterogeneity. Early works derived the $\sqrt{n}$-consistency and asymptotic normality of quantile regressors in the linear model (Bassett Jr and Koenker, 1978, 1982; Portnoy and Koenker, 1989; Pollard, 1991). Other works established statistical properties under fixed designs, where covariates are treated as deterministic (He and Shao, 1996; Koenker, 2005). More recent works have shifted toward non-asymptotic analysis with convergence rate $\mathcal{O}(1/\sqrt{n})$ under random designs, where covariates are random and prediction performance on unseen data is emphasized (Steinwart and Christmann, 2011; Catoni, 2012; Hsu et al., 2014; Loh and Wainwright, 2013; Pan and Zhou, 2021; He et al., 2023; Liu et al., 2023; Sasai and Fujisawa, 2025). Median regression is a special case of quantile regression, has also been extensively studied (Chen et al., 2008). These methods form the basis for conformalized median regression and conformalized quantile regression (Romano et al., 2019).
Efficiency analysis of conformal prediction.
Conformal prediction was developed to equip point predictions with confidence regions that provide finite-sample coverage guarantees (Papadopoulos et al., 2002; Vovk et al., 2005, 2009; Vovk, 2025). Research on its efficiency (Vovk et al., 2016; Gasparin and Ramdas, 2025) has evolved from early asymptotic convergence analyses, which established convergence rates toward the oracle prediction region (Chajewska et al., 2001; Li and Liu, 2008; Sadinle et al., 2019; Sesia and Candès, 2020; Chernozhukov et al., 2021; Izbicki et al., 2022), to generalization error-based bounds on expected set size Zecchin et al. (2024), and recently volume-minimization methods using data-driven norms (Sharma et al., 2023; Correia et al., 2024; Kiyani et al., 2024; Braun et al., 2025; Bars and Humbert, 2025; Gao et al., 2025; Srinivas, 2025).
For conditional density estimation, under $\beta$-Hölder class and $\gamma$-exponent margin conditions of the conditional density, Lei and Wasserman (2014) derived minimax-optimal rates of order $\mathcal{O}((\log m/m)^{\beta/(3\beta+1)})$ when $\gamma=1$, and showed that conditional coverage cannot generally be guaranteed in finite samples. When the quantile of $Y$ is symmetric and independent of $X$ (analogous to Assumption 4.2), Lei et al. (2018) incorporated training error into the efficiency analysis, treating $\alpha$ as a fixed constant. In contrast, our results for CQR and CMR make no assumptions on the training error and provide explicit upper bounds (B.5, C) as functions of $(n,m,\alpha)$, applicable also to adaptive prediction sets.
Under the assumptions that the quantile function of the nonconformity score is locally $\beta$-Hölder continuous, and that the worst-case empirical estimation error of the function class is bounded, Bars and Humbert (2025) derived convergence rates of the order $\mathcal{O}(m^{-\beta\kappa/2}+n^{-\beta\iota/2})$ for some $0<\iota,\kappa<1$ when the function class is finite. In the case of $\beta=1$, this rate matches our bound when $\alpha$ is treated as a fixed constant, namely $\mathcal{O}(m^{-1/2}+n^{-1/2})$. Different from analysis in Bars and Humbert (2025) that focuses on methods based on volume minimization, our work develops efficiency guarantees for CQR and CMR, without imposing assumptions on the score distribution induced by the trained model or on the estimation error. Instead, we demonstrate in the proof (especially Proposition C.2) that the required regularity conditions of the score are satisfied with high probability under mild assumptions on the underlying data distribution.
6 Experiments
This section presents evaluations of length deviation using synthetic data to access our theoretical results. Real-world experiments are deferred to Appendix E due to space constraints.
Experiment setup. The data generation procedure is described in Appendix D.1. All experiments employ linear models trained with SGD for one epoch using a batch size of $64$. Learning rates are selected via successive halving over the range $10^{-5},1]$. We evaluate miscoverage levels $\alpha\in\{0.01,0.025,0.05,0.075,0.1,0.125,0.15,0.175,0.2\}$. Reported results are averaged over $20$ independent trials, and length deviations are computed on $2000$ test samples.
We denote the expected length deviation as $\Delta$. We empirically assess the upper bound of $\Delta$ in Theorem 3.2, of order $\mathcal{O}(\frac{1}{\sqrt{n}}+\frac{1}{n\alpha^{2}}+\frac{1}{\sqrt{m}}+\exp(-\alpha^{2}m))$ from three perspectives.
  • Effect of training size $n$. With a large calibration set ($m=5000$), the calibration error is negligible, and the theoretical bound simplifies to $\mathcal{O}(1/\sqrt{n}+1/(n\alpha^{2}))$. The theory predicts that a linear regression of $\log\Delta$ on $\log n$, i.e.,
    $\log\Delta\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{1}}\log n+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{2}},$ (23)
    yields a slope $a_{1}$ that transitions from 1-1 to 1/2-{1}/{2} as $\alpha$ increases. We confirm this trend empirically. For each $\alpha$, we train models over $n$ ranging from $200$ to $20000$ (Fig. 3a) and fit the regression model (23) (the inset in Fig. 3a shows an example) to obtain slope $a_{1}$ and intercept $a_{2})$. The resulting $(\alpha,a_{1})$ pairs, shown by the red curve in Fig. 3c, validate that the slope shifts from approximately 1-1 to 1/2-{1}/{2} as $\alpha$ grows, reflecting the transition of the dominant term in the bound from $\mathcal{O}(1/(n\alpha^{2}))$ to $\mathcal{O}(1/\sqrt{n})$. The intercept $a_{2}$ depends on $\log\alpha$, as discussed below.
  • Effect of miscoverage level $\alpha$. In the regime where $(n\alpha^{2})^{-1}$ dominates, $\Delta$ is expected to follow a power-law scaling of order $\alpha^{-2}$. To examine this, we further regress the fitted intercepts $a_{2}$ in (23) on $\log\alpha$:
    ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{2}}\sim b_{1}\log\alpha+b_{2}.$
    Together with (23), the estimated coefficient $b_{1}=-2.24$ (Fig. 3d) implies that $\Delta\sim\alpha^{-2.24}$. This aligns with the theoretical upper bound of order $\mathcal{O}(\alpha^{-2})$. Appendix D.2 provides an additional verification for the existence of this regime.
  • Effect of calibration size $m$. Using the ground-truth parameter $\theta^{*}$, we vary the calibration set size $m$ ranging from $100$ to $3000$, ensuring that the resulting length deviation depends only on $m$ and $\alpha$. As illustrated in Fig. 3b, the deviation decreases consistently with larger calibration sets. On a log–log scale, the slope approximately approaches 0.5-0.5, reflecting the increasing dominance of the $\mathcal{O}({1}/{\sqrt{m}})$ term in the bound. Meanwhile, the exponential term $\exp(-\alpha^{2}m)$ decays quickly for modest values of $m$ and becomes negligible thereafter.
Figure 3: The length deviation of conformalized quantile regression in synthetic data experiments.
7 Conclusion
This paper studies the efficiency of conformalized quantile regression (CQR) and conformalized median regression (CMR) through the lens of the expected length deviation, defined as the discrepancy between the coverage-guaranteed prediction set size and the oracle interval length. Our analysis explicitly accounts for randomness introduced by training, finite-sample calibration, and test evaluation. Under mild assumptions on the data distribution, we provide, to the best of our knowledge, the first non-asymptotic convergence rate of the form: $\mathcal{O}(n^{-1/2}+n^{-1}\alpha^{-2}+m^{-1/2}+\exp(-\alpha^{2}m))$, which highlights a fine-grained effect of the miscoverage level $\alpha$. Empirical results closely align with the theoretical findings.
Appendix A Discussion on Upper bounds for Quantile Estimation Error
In Theorem 3.1, we provide the upper bound $\mathcal{O}(n^{-1})$ for the quantile estimation error using standard SGD when the objective (4) is strongly convex under Assumption 3.2 and 3.3. We notice that Assumption 3.2 can be relaxed as discussed in Remark 3.2.
We also note that variance-reduction techniques, such as SAG (Schmidt et al., 2017), SVRG (Johnson and Zhang, 2013), and SAGA (Defazio et al., 2014), achieve a linear convergence rate under strong convexity, meaning that the error decays geometrically in the number of iterations $\mathcal{O}(\exp(-cn))$. This is in sharp contrast to the $\mathcal{O}(n^{-1})$ rate of standard SGD. Since our focus in this paper is not on improving the convergence rate of quantile estimation, we present results under the standard SGD rate in Theorem 3.1. Nevertheless, accelerated (linear) rates can be obtained by incorporating variance-reduction techniques following the same proof strategy developed here.
Appendix B Proofs of Results in CQR
To proceed, we first define some notations as follows.
$\mathcal{E}_{\gamma}\left(X,\theta_{n}\left(\gamma\right)\right)$ $:=\left|t_{\gamma}\left(X;\theta_{n}\left(\gamma\right)\right)-t_{\gamma}\left(X;\theta^{*}\left(\gamma\right)\right)\right|\geq 0$ (24)
$\Delta\left(X,\vartheta_{n}\right)$ $:=\max\left\{\mathcal{E}_{\alpha/2}\left(X,\underline{\theta}_{n}\right),\;\mathcal{E}_{1-\alpha/2}\left(X,\bar{\theta}_{n}\right)\right\}\geq 0$ (25)
$S^{*}\left(X,Y\right)$ $:=\max\left\{t_{\alpha/2}\left(X;\underline{\theta}^{*}\right)-Y,\;Y-t_{1-\alpha/2}\left(X;\bar{\theta}^{*}\right)\right\}$ (26)
$=\max\left\{q_{\alpha/2}\left(Y\mid X\right)-Y,\;Y-q_{1-\alpha/2}\left(Y\mid X\right)\right\}$
$M\left(\vartheta_{n}\right)$ $:=\max\left\{\left\|\left(\underline{\theta}_{n}-\underline{\theta}^{*}\right)\right\|_{2},\;\left\|\left(\bar{\theta}_{n}-\bar{\theta}^{*}\right)\right\|_{2}\right\}$ (27)
Let $\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}$ denote the empirical c.d.f. from $m$ i.i.d. calibration scores given $\vartheta_{n}$, i.e.,
$\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(s\right)=\frac{1}{m}\sum_{j=1}^{m}\mathbbm{1}\{S_{j}\leq s\},\qquad S_{j}\overset{\text{i.i.d.}}{\sim}S\mid\vartheta_{n}$
B.1 Proof of Theorem 3.1
See 3.1
To prove Theorem 3.1, we first show that $\ell_{\gamma}\left(\theta\right)$ in (4) is strongly convex and smooth with respect to $\theta^{*}(\gamma)$, as stated below in Proposition B.1. The proof of Proposition B.1 further relies on Lemma B.2 and Lemma B.3 for the gradient and the Hessian of $\ell_{\gamma}\left(\theta\right)$.
Proposition B.1.
Under Assumption 3.3, and if $\mathbb{E}\left[\|X\|^{2}\right]<\infty$, the objective $\ell_{\gamma}\left(\theta\right)$ in (4) satisfies
$\frac{f_{\min}}{2}\|\theta-\theta^{*}\left(\gamma\right)\|_{\Sigma}^{2}\leq\ell_{\gamma}\left(\theta\right)-\ell_{\gamma}\left(\theta^{*}\left(\gamma\right)\right)\leq\frac{f_{\max}}{2}\|\theta-\theta^{*}\left(\gamma\right)\|_{\Sigma}^{2}$ (28)
If Assumption 3.2 furthermore holds, then
$\frac{f_{\min}\lambda_{\min}}{2}\|\theta-\theta^{*}\left(\gamma\right)\|_{2}^{2}\leq\ell_{\gamma}\left(\theta\right)-\ell_{\gamma}\left(\theta^{*}\left(\gamma\right)\right)\leq\frac{f_{\max}\lambda_{\max}}{2}\|\theta-\theta^{*}\left(\gamma\right)\|_{2}^{2}$ (29)
Proof.
To prove this proposition, we first need Lemma B.2 and Lemma B.3 to calculate the gradient and the Hessian of $\ell_{\gamma}\left(\theta\right)$. By Lemma B.2,
$\nabla\ell_{\gamma}\left(\theta^{*}\left(\gamma\right)\right)$ $=\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(\left(\theta^{*}\left(\gamma\right)\right)^{\top}X\mid X\right)-\gamma\right)X\right]$
$=\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(q_{\gamma}\left(Y\mid X\right)\right)-\gamma\right)X\right]$
$=0$
By Lemma B.3, $\nabla^{2}\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[f_{Y\mid X}\left(\theta^{\top}X\mid X\right)XX^{\top}\right]$. By Assumption 3.3, $\forall v\in\mathbb{R}^{d}$,
$f_{\min}\|v\|_{\Sigma}^{2}=f_{\min}\mathbb{E}_{X}\left[\left(X^{\top}v\right)^{2}\right]$ $\leq\mathbb{E}_{X}\left[f_{Y\mid X}\left(\theta^{\top}X\mid X\right)\left(X^{\top}v\right)^{2}\right]$
$\leq f_{\max}\mathbb{E}_{X}\left[\left(X^{\top}v\right)^{2}\right]=f_{\max}\|v\|_{\Sigma}^{2}$
Hence, $f_{\min}\Sigma\preceq\nabla^{2}\ell_{\gamma}\left(\theta\right)\preceq f_{\max}\Sigma$ for any $\theta\in\Theta$. By Taylor’s Formula,
$\ell_{\gamma}\left(\theta\right)-\ell_{\gamma}\left(\theta^{*}\left(\gamma\right)\right)=\int_{0}^{1}\left(1-u\right)\left(\theta-\theta^{*}\left(\gamma\right)\right)^{\top}\nabla^{2}\ell_{\gamma}\left(\theta^{*}+u\left(\theta-\theta^{*}\left(\gamma\right)\right)\right)\left(\theta-\theta^{*}\left(\gamma\right)\right)du$
Since
$f_{\min}\|\theta-\theta^{*}\left(\gamma\right)\|_{\Sigma}$ $\leq\left(\theta-\theta^{*}\left(\gamma\right)\right)^{\top}\nabla^{2}\ell_{\gamma}\left(\theta^{*}+u\left(\theta-\theta^{*}\left(\gamma\right)\right)\right)\left(\theta-\theta^{*}\left(\gamma\right)\right)$
$\leq f_{\max}\|\theta-\theta^{*}\left(\gamma\right)\|_{\Sigma}$
and $\int_{0}^{1}\left(1-u\right)\;du=1/2$, we have
$\frac{f_{\min}}{2}\|\theta-\theta^{*}\left(\gamma\right)\|_{\Sigma}^{2}\leq\ell_{\gamma}\left(\theta\right)-\ell_{\gamma}\left(\theta^{*}\left(\gamma\right)\right)\leq\frac{f_{\max}}{2}\|\theta-\theta^{*}\left(\gamma\right)\|_{\Sigma}^{2}$
Lemma B.2.
Suppose (11) in Assumption 3.3 is true, if $\mathbb{E}\left[\|X\|_{2}\right]<\infty$, then
$\nabla\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X,Y}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)X\right]=\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(\theta^{\top}X\mid X\right)-\gamma\right)X\right]$ (30)
Proof.
The key idea is to show that the interchange of differentiation and expectation is valid according to the dominated convergence theorem. For $\theta\in\Theta$, it holds that
$\mathbb{P}\left[Y=\theta^{\top}X\right]$ $=\mathbb{E}_{\left(X,Y\right)}\left[\mathbbm{1}\left\{Y=\theta^{\top}X\right\}\right]$
$=\mathbb{E}_{X}\left[\mathbb{E}_{Y\mid X}\left[\mathbbm{1}\left\{Y=\theta^{\top}X\right\}\mid X\right]\right]$
$=\mathbb{E}_{X}\left[\mathbb{P}\left[Y=\theta^{\top}X\mid X\right]\right]$
Since (11) in Assumption 3.3 is true, the p.d.f $f_{Y\mid X}\left(Y\mid X\right)$ exists for each $x\in\mathcal{X}$. Thus,
$\mathbb{P}\left[Y=\theta^{\top}x\mid X=x\right]=\int_{\{\theta^{\top}x\}}f_{Y\mid X}\left(Y\mid X\right)dy=0.$
Thus, $\mathbb{P}\left[Y=t_{\gamma}\left(X;\theta\right)\right]=\mathbb{P}\left[Y=\theta^{\top}X\right]=\mathbb{E}[0]=0$.
For $\left(x,y\right)\in\mathcal{X}\times\mathcal{Y}$, if $y\neq t_{\gamma}\left(x;\theta\right)$, the directional derivative of $L_{\gamma}\left(\theta^{\top}x,y\right)$ at $\theta$ along vector $v$ is
$D_{v}L_{\gamma}\left(\theta^{\top}x,y\right)$ $=\lim_{\rho\to 0}\frac{L_{\gamma}\left(\left(\theta+\rho v\right)^{\top}x,y\right)-L_{\gamma}\left(\theta^{\top}x,y\right)}{\|v\|_{2}\rho}$
$=\frac{1}{\|v\|}\left.\frac{d}{d\rho}L_{\gamma}\left(\left(\theta+\rho v\right)^{\top}x,y\right)\right|_{\rho=0}$
$=\left(\mathbbm{1}\left\{y<\theta^{\top}x\right\}-\gamma\right)x^{\top}\frac{v}{\|v\|}$
Moreover, since $L_{\gamma}\left(t,y\right)$ is $1$-Lipschitz with respect to $t$,
$\left|\frac{L_{\gamma}\left(\left(\theta+\rho v\right)^{\top}x,y\right)-L_{\gamma}\left(\theta^{\top}x,y\right)}{\|v\|_{2}\rho}\right|$ $=\frac{1}{\|v\|_{2}\rho}\left|L_{\gamma}\left(\left(\theta+\rho v\right)^{\top}x,y\right)-L_{\gamma}\left(\theta^{\top}x,y\right)\right|$
$\leq\frac{1}{\|v\|_{2}\rho}\|\left(\theta+\rho v\right)^{\top}x-\theta^{\top}x\|_{2}$
$\leq\|x\|$
Since we assume $\mathbb{E}\left[\|X\|_{2}\right]<\infty$, by the dominated convergence theorem,
$D_{v}\ell_{\gamma}\left(\theta\right)$ $=D_{v}\mathbb{E}_{X,Y}\left[L_{\gamma}\left(\theta^{\top}X,Y\right)\right]$
$=\lim_{\rho\to 0}\frac{\mathbb{E}_{X,Y}\left[L_{\gamma}\left(\left(\theta+\rho v\right)^{\top}X,Y\right)\right]-\mathbb{E}_{X,Y}\left[L_{\gamma}\left(\theta^{\top}X,Y\right)\right]}{\|v\|_{2}\rho}$
$=\lim_{\rho\to 0}\mathbb{E}_{X,Y}\left[\frac{L_{\gamma}\left(\left(\theta+\rho v\right)^{\top}X,Y\right)-L_{\gamma}\left(\theta^{\top}X,Y\right)}{\|v\|_{2}\rho}\right]$
$=\mathbb{E}_{X,Y}\left[\lim_{\rho\to 0}\frac{L_{\gamma}\left(\left(\theta+\rho v\right)^{\top}X,Y\right)-L_{\gamma}\left(\theta^{\top}X,Y\right)}{\|v\|_{2}\rho}\right]$
$=\mathbb{E}_{X,Y}\left[D_{v}L_{\gamma}\left(\theta^{\top}X,Y\right)\right]$
$=\mathbb{E}_{X,Y}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)X\right]^{\top}\frac{v}{\|v\|}$
Hence,
$\nabla\ell_{\gamma}\left(\theta\right)$ $=\mathbb{E}_{X,Y}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)X\right]$
$=\mathbb{E}_{X}\left[\mathbb{E}_{Y\mid X}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)X\mid X\right]\right]$
$=\mathbb{E}_{X}\left[\mathbb{E}_{Y\mid X}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)\mid X\right]X\right]$
$=\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(\theta^{\top}X\mid X\right)-\gamma\right)X\right]$
Lemma B.3.
Under Assumption 3.3, if $\mathbb{E}\left[\|X\|^{2}\right]<\infty$, then
$\nabla^{2}\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[f_{Y\mid X}\left(\theta^{\top}X\mid X\right)XX^{\top}\right]$ (31)
Proof.
By Assumption, $\mathbb{E}\left[\|X\|_{2}\right]\leq\sqrt{\mathbb{E}\left[\|X\|^{2}\right]}<\infty$. Then, by Lemma B.2,
$\nabla\ell_{\gamma}\left(\theta\right)$ $=\mathbb{E}_{X,Y}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)X\right]$
$=\mathbb{E}_{X}\left[\mathbb{E}_{Y\mid X}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)X\mid X\right]\right]$
$=\mathbb{E}_{X}\left[\mathbb{E}_{Y\mid X}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}-\gamma\right)\mid X\right]X\right]$
$=\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(\theta^{\top}X\mid X\right)-\gamma\right)X\right]$
To prove the lemma, the key point is to show that the interchange of differentiation and expectation is valid, as in the proof of Lemma B.2.
$\lim_{\rho\to 0}\frac{\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-\gamma\right)x-\left(F_{Y\mid X}\left(\theta^{\top}x\mid X\right)-\gamma\right)x}{\|v\|_{2}\rho}$
$=\lim_{\rho\to 0}\frac{1}{\|v\|_{2}\rho}\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-F_{Y\mid X}\left(\theta^{\top}x\mid X\right)\right)x$
$=x\cdot\frac{v^{\top}x}{\|v\|}\lim_{\rho\to 0}\frac{1}{\rho v^{\top}x}\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-F_{Y\mid X}\left(\theta^{\top}x\mid X\right)\right)$
According to the mean value theorem, there exists $\xi\left(x\right)$ in $\left(\theta^{\top}x,\theta^{\top}x+\rho v^{\top}x\right)$ such that
$\frac{1}{\rho v^{\top}x}\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-F_{Y\mid X}\left(\theta^{\top}X\mid X\right)\right)=f_{Y\mid X}\left(\xi\left(x\right)\mid X\right)$
Hence,
$\lim_{\rho\to 0}\frac{1}{\rho v^{\top}x}\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-F_{Y\mid X}\left(\theta^{\top}X\mid X\right)\right)=\lim_{\rho\to 0}f_{Y\mid X}\left(\xi\left(x\right)\mid X\right)$
Since $f_{Y\mid X}\left(Y\mid X\right)$ is continuous for $\mathcal{P}_{X}$-almost every $x\in\mathcal{X}$, we have for $\mathcal{P}_{X}$-almost every $x\in\mathcal{X}$,
$\lim_{\rho\to 0}f_{Y\mid X}\left(\xi\left(x\right)\mid X\right)=f_{Y\mid X}\left(\theta^{\top}X\mid X\right)$
Hence, for $\mathcal{P}_{X}$-almost every $x\in\mathcal{X}$,
$\lim_{\rho\to 0}\frac{\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-\gamma\right)x-\left(F_{Y\mid X}\left(\theta^{\top}X\mid X\right)-\gamma\right)x}{\|v\|_{2}\rho}=f_{Y\mid X}\left(\theta^{\top}X\mid X\right)\frac{xx^{\top}v}{\|v\|}$
Since (11) in Assumption 3.3 is true, for any $x\in\mathcal{X}$, $F_{Y\mid X}$ is $f_{\max}$-Lipschitz.
$\left|\frac{\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-\gamma\right)x-\left(F_{Y\mid X}\left(\theta^{\top}X\mid X\right)-\gamma\right)x}{\|v\|_{2}\rho}\right|$
$=\frac{1}{\|v\|_{2}\rho}\left|\left(F_{Y\mid X}\left(\theta^{\top}x+\rho v^{\top}x\mid X\right)-F_{Y\mid X}\left(\theta^{\top}X\mid X\right)\right)\right|\|x\|_{2}$
$\leq\frac{1}{\|v\|_{2}\rho}f_{\max}\rho\|v\|_{2}\|x\|^{2}=f_{\max}\|x\|^{2}$
Since $\mathbb{E}\left[\|X\|^{2}\right]<\infty$, it holds that $\mathbb{E}\left[f_{\max}\|X\|^{2}\right]<\infty$. Therefore, by the dominated convergence theorem, the directional derivative of $\nabla\ell_{\gamma}\left(\theta\right)$ at $\theta$ along vector $v$ is
$D_{v}\left(\nabla\ell_{\gamma}\left(\theta\right)\right)$
$=D_{v}\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(\theta^{\top}X\mid X\right)-\gamma\right)X\right]$
$=\lim_{\rho\to 0}\frac{\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(\theta^{\top}X+\rho v^{\top}X\mid X\right)-\gamma\right)X\right]-\mathbb{E}_{X}\left[\left(F_{Y\mid X}\left(\theta^{\top}X\mid X\right)-\gamma\right)X\right]}{\|v\|_{2}\rho}$
$=\lim_{\rho\to 0}\mathbb{E}_{X}\left[\frac{1}{\|v\|_{2}\rho}\left(F_{Y\mid X}\left(\theta^{\top}X+\rho v^{\top}X\mid X\right)-F_{Y\mid X}\left(\theta^{\top}X\mid X\right)\right)X\right]$
$=\mathbb{E}_{X}\left[\lim_{\rho\to 0}\frac{1}{\|v\|_{2}\rho}\left(F_{Y\mid X}\left(\theta^{\top}X+\rho v^{\top}X\mid X\right)-F_{Y\mid X}\left(\theta^{\top}X\mid X\right)\right)X\right]$
$=\mathbb{E}_{X}\left[f_{Y\mid X}\left(\theta^{\top}X\mid X\right)XX^{\top}\right]\frac{v}{\|v\|}$
Hence, $\nabla^{2}\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[f_{Y\mid X}\left(\theta^{\top}X\mid X\right)XX^{\top}\right]$. ∎
With Proposition B.1, we are ready to apply Theorem B.4 for SGD and get Corollary B.1.
Theorem B.4 (Section 3 in Rakhlin et al. (2012)).
Suppose the loss function $\ell$ is $\lambda$-strongly convex and $\mu$-smooth with respect to a minimizer $\theta^{*}$ over $\Theta$, and $\mathbb{E}[\|g_{t}\|^{2}]\leq G^{2}$. Then taking $\eta_{t}=1/\lambda t$, it holds for any $n$ that
$\mathbb{E}_{\theta_{n}}\left[f\left(\theta_{n}\right)-f\left(\theta^{*}\right)\right]\leq\frac{2\mu G^{2}}{\lambda^{2}n}.$ (32)
Corollary B.1 (Upper Bound of Extra Loss).
Suppose Assumption 3.1, 3.2 and 3.3 hold. Let $\mathcal{D}_{train}:=\{\left(X_{i},Y_{i}\right)\}_{i=1}^{n}$ be the set of training samples and $\theta_{n}$ be the estimator by optimizing stochastic pinball loss (4) produced by SGD (7). Taking $\eta_{t}=1/\left(\lambda_{\min}f_{\min}t\right)$, it holds that
$\mathbb{E}_{\theta_{n}}\left[\ell_{\gamma}\left(\theta_{n}\left(\gamma\right)\right)-\ell_{\gamma}\left(\theta^{*}\left(\gamma\right)\right)\right]\leq\frac{2\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{2}f_{\min}^{2}n}.$ (33)
Proof.
In this proof, we denote $\theta_{n}\left(\gamma\right)$ by $\theta_{n}$ for simplicity. By Lemma B.2, $\nabla\ell_{\gamma}\left(\theta\right)=\mathbb{E}_{X}\left[\left(\mathbbm{1}\left\{Y<\theta^{\top}X\right\}\right)X\right]$. Then,
$\mathbb{E}_{X,\theta_{n}}\left[\left\|\nabla\ell_{\gamma}\left(\theta_{n}\right)\right\|^{2}\right]$ $=\mathbb{E}_{\theta_{n}}\left[\left\|\mathbb{E}_{X}\left[\left(\mathbbm{1}\left\{Y<\theta_{n}^{\top}X\right\}\right)X\right]\right\|^{2}\right]$
$=\mathbb{E}_{\theta_{n}}\left[\mathbb{E}_{X}\left[\left\|\left(\mathbbm{1}\left\{Y<\theta_{n}^{\top}X\right\}\right)X\right\|\right]^{2}\right]$
$\leq\mathbb{E}_{X}\left[\left\|X\right\|\right]^{2}\leq\lambda_{\max}d$
where the last inequality is by Assumption 3.2,
$\mathbb{E}\left[\left\|X\right\|\right]^{2}\leq\mathbb{E}\left[\|X\|^{2}\right]=\mathbb{E}\left[\text{trace}\left(XX^{\top}\right)\right]=\text{trace}\left(\mathbb{E}\left[XX^{\top}\right]\right)\leq\text{trace}\left(\lambda_{\max}I\right)=d\lambda_{\max}$
The corollary then follows from Proposition B.1 and Theorem B.4. ∎
Now we are ready to prove Theorem 3.1. In this proof, we denote $\theta_{n}\left(\gamma\right),\theta^{*}\left(\gamma\right)$ by $\theta_{n},\theta^{*}$, respectively, for simplicity. By Proposition B.1,
$\|\theta_{n}-\theta^{*}\|_{\Sigma}^{2}$ $\leq\frac{2}{f_{\min}}\left(\ell\left(\theta_{n}\right)-\ell\left(\theta^{*}\right)\right)$
$\|\theta_{n}-\theta^{*}\|_{2}^{2}$ $\leq\frac{2}{f_{\min}\lambda_{\min}}\left(\ell\left(\theta_{n}\right)-\ell\left(\theta^{*}\right)\right)$
Since the test sample $\left(X,Y\right)$ is sampled independently of the set of the training samples $\{\left(X_{i},Y_{i}\right)\}_{i=1}^{n}$, and $\theta_{n}$ is a function of $\{\left(X_{i},Y_{i}\right)\}_{i=1}^{n}$, $\theta_{n}$ is independent of $X$.
$\mathbb{E}_{\theta_{n},X}\left[\left(t\left(X;\theta_{n}\right)-t\left(X;\theta^{*}\right)\right)^{2}\right]$ $=\mathbb{E}_{\theta_{n},X}\left[\left(\left(\theta_{n}-\theta^{*}\right)^{\top}X\right)^{2}\right]$
$=\mathbb{E}_{\theta_{n}}\left[\mathbb{E}_{X}\left[\left(\theta_{n}-\theta^{*}\right)^{\top}XX^{\top}\left(\theta_{n}-\theta^{*}\right)|\theta_{n}\right]\right]$
$=\mathbb{E}_{\theta_{n}}\left[\left(\theta_{n}-\theta^{*}\right)^{\top}\mathbb{E}_{X}\left[XX^{\top}\right]\left(\theta_{n}-\theta^{*}\right)\right]$
$=\mathbb{E}_{\theta_{n}}\left[\left\|\theta_{n}-\theta^{*}\right\|_{\Sigma}^{2}\right]$
Hence, by Corollary B.1, $\mathbb{E}_{\theta_{n}}[\|\theta_{n}-\theta^{*}\|_{\Sigma}^{2}]\leq\frac{2}{f_{\min}}\mathbb{E}_{\theta_{n}}[\left(\ell\left(\theta_{n}\right)-\ell\left(\theta^{*}\right)\right)]\leq\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{3}f_{\min}^{2}n}$.
This completes the proof of Theorem 3.1.
B.2 Proof of Proposition B.5
Proposition B.5 (Population quantile of the score).
In CQR, if $F_{Y\mid X}\left(Y\mid X=x\right)$ is continuous for all $x\in\mathcal{X}$, then
$\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|$ $\leq B\max\left\{\left\|\underline{\theta}_{n}-\underline{\theta}^{*}\right\|_{2},\left\|\bar{\theta}_{n}-\bar{\theta}^{*}\right\|_{2}\right\}$ (34)
Suppose Assumptions 3.13.3 hold,
$\mathbb{E}_{\vartheta_{n}}\left[\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\right]$ $\leq\frac{2B\ \lambda_{\max}\sqrt{2f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\sqrt{\frac{1}{n}}$ (35)
The proof of Proposition B.5 relies on the following lemma.
Lemma B.6.
Suppose $F_{Y\mid X}\left(Y\mid X\right)$ is continuous for each $x\in\mathcal{X}$. Then,
$\left|q_{1-\alpha}\left(S\mid X,\vartheta_{n}\right)\right|\leq\Delta\left(X,\vartheta_{n}\right)$ (36)
where $q_{1-\alpha}\left(S\mid X,\vartheta_{n}\right)$ denotes the $\left(1-\alpha\right)$-quantile of $S$ given $X,\check{\theta}_{n}$.
Proof.
By the definitions (24, 25, 26),
$S\left(X,Y;\vartheta_{n}\right)$ $:=\max\left\{t_{\alpha/2}\left(X;\underline{\theta}_{n}\right)-Y,\;Y-t_{1-\alpha/2}\left(X;\bar{\theta}_{n}\right)\right\}$
$\leq\max\left\{\mathcal{E}_{\alpha/2}\left(X,\underline{\theta}_{n}\right)+q_{\alpha/2}\left(Y\mid X\right)-Y,\;\mathcal{E}_{1-\alpha/2}\left(X,\bar{\theta}_{n}\right)+Y-q_{1-\alpha/2}\left(Y\mid X\right)\right\}$
$\leq\Delta\left(X,\vartheta_{n}\right)+S^{*}\left(X,Y\right)$ (37)
where the last inequality is because $\max\{u_{1}+v_{1},u_{2}+v_{2}\}\leq\max\{u_{1},u_{2}\}+\max\{v_{1},v_{2}\}$.
Similarly,
$S\left(X,Y;\vartheta_{n}\right)$ $:=\max\left\{t_{\alpha/2}\left(X;\underline{\theta}_{n}\right)-Y,\;Y-t_{1-\alpha/2}\left(X;\bar{\theta}_{n}\right)\right\}$
$\geq\max\left\{q_{\alpha/2}\left(Y\mid X\right)-Y-\mathcal{E}_{\alpha/2}\left(X,\underline{\theta}_{n}\right),\;Y-q_{1-\alpha/2}\left(Y\mid X\right)-\mathcal{E}_{1-\alpha/2}\left(X,\bar{\theta}_{n}\right)\right\}$
$=S^{*}\left(X,Y\right)-\Delta\left(X,\vartheta_{n}\right)$ (38)
where the last inequality is because $\max\{u_{1}-v_{1},u_{2}-v_{2}\}\geq\max\{u_{1},u_{2}\}-\max\{v_{1},v_{2}\}$.
Note that $S^{*}\left(X,Y\right)\leq 0$ is equivalent to $q_{\alpha/2}\left(Y\mid X\right)\leq Y\leq q_{1-\alpha/2}\left(Y\mid X\right)$. Since $F_{Y\mid X}$ is continuous,
$\mathbb{P}\left[q_{\alpha/2}\left(Y\mid X\right)\leq Y\leq q_{1-\alpha/2}\left(Y\mid X\right)\mid X\right]=1-\alpha$
Hence, $\mathbb{P}[S^{*}\left(X,Y\right)\leq 0|X]=1-\alpha$. Let $q_{1-\alpha}\left(S^{*}\mid X\right)$ be the $\left(1-\alpha\right)$-quantile of $S^{*}$ given $X$. Since $X$ is given, and $F_{Y\mid X}$ is continuous, $F_{S^{*}|X}$ is continuous. Then, $q_{1-\alpha}\left(S^{*}\mid X\right)=0$. Conditional on $X,\vartheta_{n}$, $\Delta\left(X,\vartheta_{n}\right)$ is deterministic. By (37), we have
$\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\leq u\mid X,\vartheta_{n}\right]\geq\mathbb{P}\left[\Delta\left(X,\vartheta_{n}\right)+S^{*}\left(X,Y\right)\leq u\mid X,\vartheta_{n}\right]$
$\Longrightarrow\;\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\leq\Delta\left(X,\vartheta_{n}\right)\mid X,\vartheta_{n}\right]\geq\mathbb{P}\left[S^{*}\left(X,Y\right)\leq 0\mid X\right]=1-\alpha$
Then, $q_{1-\alpha}\left(S\mid X,\vartheta_{n}\right)\leq\Delta\left(X,\vartheta_{n}\right)$. By (38), we have
$\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\leq u\mid X,\vartheta_{n}\right]\leq\mathbb{P}\left[S^{*}\left(X,Y\right)-\Delta\left(X,\vartheta_{n}\right)\leq u\mid X,\vartheta_{n}\right]$
$\Longrightarrow\;\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\leq-\Delta\left(X,\vartheta_{n}\right)\mid X,\vartheta_{n}\right]\leq\mathbb{P}\left[S^{*}\left(X,Y\right)\leq 0\mid X\right]=1-\alpha$
Then, $q_{1-\alpha}\left(S\mid X,\vartheta_{n}\right)\geq-\Delta\left(X,\vartheta_{n}\right)$. ∎
For $\gamma\in\{\frac{\alpha}{2},1-\frac{\alpha}{2}\}$,
$\mathcal{E}_{\gamma}\left(X,\theta_{n}\left(\gamma\right)\right)=\left|\left(\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\right)^{\top}X\right|\leq\left\|\left(\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\right)\right\|_{2}\left\|X\right\|_{2}\leq B\left\|\left(\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\right)\right\|_{2}$
where the last inequality is from the fact that the norm of $x\in\mathcal{X}$ is bounded by $B$. Then,
$\Delta\left(X,\vartheta_{n}\right)\leq B\max\left\{\left\|\left(\underline{\theta}_{n}-\underline{\theta}^{*}\right)\right\|_{2},\;\left\|\left(\bar{\theta}_{n}-\bar{\theta}^{*}\right)\right\|_{2}\right\}=B\cdot M\left(\vartheta_{n}\right)$
By Lemma B.6, $\left|q_{1-\alpha}\left(S\mid X,\vartheta_{n}\right)\right|\leq\Delta\left(X,\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)$. Then,
$\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)\mid X,\vartheta_{n}\right]\geq 1-\alpha$
$\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\geq-B\cdot M\left(\vartheta_{n}\right)\mid X,\vartheta_{n}\right]\leq 1-\alpha$
Then, removing the conditioning on $X$,
$\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)\mid\vartheta_{n}\right]$
$=\mathbb{E}_{X,Y|\vartheta_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)\right\}\mid\vartheta_{n}\right]$
$=\mathbb{E}_{X|\vartheta_{n}}\left[\mathbb{E}_{Y|X,\vartheta_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)\right\}\mid X,\vartheta_{n}\right]\mid\vartheta_{n}\right]$
$=\mathbb{E}_{X|\vartheta_{n}}\left[\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)\mid X,\vartheta_{n}\right]\mid\vartheta_{n}\right]$
$\geq\mathbb{E}_{X|\vartheta_{n}}\left[1-\alpha\mid\vartheta_{n}\right]=1-\alpha$
Hence, $q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\leq B\cdot M\left(\vartheta_{n}\right)$. And by similar arguments as below, $q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\geq-B\cdot M\left(\vartheta_{n}\right)$.
$\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\geq-B\cdot M\left(\vartheta_{n}\right)\mid\vartheta_{n}\right]$
$=\mathbb{E}_{X,Y|\vartheta_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\vartheta_{n}\right)\geq-B\cdot M\left(\vartheta_{n}\right)\right\}\mid\vartheta_{n}\right]$
$=\mathbb{E}_{X|\vartheta_{n}}\left[\mathbb{E}_{Y|X,\vartheta_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\vartheta_{n}\right)\geq-B\cdot M\left(\vartheta_{n}\right)\right\}\mid X,\vartheta_{n}\right]\mid\vartheta_{n}\right]$
$=\mathbb{E}_{X|\vartheta_{n}}\left[\mathbb{P}\left[S\left(X,Y;\vartheta_{n}\right)\geq-B\cdot M\left(\vartheta_{n}\right)\mid X,\vartheta_{n}\right]\mid\vartheta_{n}\right]$
$\leq\mathbb{E}_{X|\vartheta_{n}}\left[1-\alpha\mid\vartheta_{n}\right]=1-\alpha$
Therefore, $\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\leq B\cdot M\left(\vartheta_{n}\right)$. Then,
$\mathbb{E}_{\vartheta_{n}}\left[\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\right]$ $\leq B\ \mathbb{E}_{\vartheta_{n}}\left[M\left(\vartheta_{n}\right)\right]$
$\leq B\ \mathbb{E}_{\vartheta_{n}}\left[\sqrt{\left\|\left(\underline{\theta}_{n}-\underline{\theta}^{*}\right)\right\|_{2}^{2}+\left\|\left(\bar{\theta}_{n}-\bar{\theta}^{*}\right)\right\|_{2}^{2}}\right]$
$\leq B\ \sqrt{\mathbb{E}_{\vartheta_{n}}\left[\left\|\left(\underline{\theta}_{n}-\underline{\theta}^{*}\right)\right\|_{2}^{2}+\left\|\left(\bar{\theta}_{n}-\bar{\theta}^{*}\right)\right\|_{2}^{2}\right]}$
$\leq B\ \sqrt{\mathbb{E}_{\vartheta_{n}}\left[\left\|\left(\underline{\theta}_{n}-\underline{\theta}^{*}\right)\right\|_{2}^{2}\right]+\mathbb{E}_{\vartheta_{n}}\left[\left\|\left(\bar{\theta}_{n}-\bar{\theta}^{*}\right)\right\|_{2}^{2}\right]}$
$\leq B\ \sqrt{\frac{8\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{4}f_{\min}^{2}n}}=\frac{2B\ \lambda_{\max}\sqrt{2f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\sqrt{\frac{1}{n}}$
where the second inequality is from $\max\{a,b\}\leq\sqrt{a^{2}+b^{2}}$, the third inequality is by Jensen’s inequality, and the last inequality is from Theorem 3.1.
This completes the proof of Proposition B.5.
B.3 Proof of Proposition B.7
Proposition B.7 (Population finite-sample score-quantile gap).
In CQR, Suppose Assumptions 3.13.3 hold, if $m>8H/\min\{\alpha,1-\alpha\}$ for $H$ in (12), then
$\mathbb{E}_{\vartheta_{n}}\left[|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|\right]\leq\frac{1}{f_{\min}m}+\frac{1056Rf_{\max}^{3}\lambda_{\max}^{2}B^{2}d}{\min\{\alpha^{2},(1-\alpha)^{2}\}\lambda_{\min}^{4}f_{\min}^{2}n}$
To prove Proposition B.7, we first need the following critical proposition:
Proposition B.8.
Suppose $\alpha\in\left(0,1\right)$ is a constant. Define
$\beta:=\min\left\{\frac{\alpha}{2f_{\max}},\ \frac{1-\alpha}{2f_{\max}}\right\}\qquad A:=\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{4}f_{\min}^{2}}\qquad\varepsilon_{n}:=B\sqrt{\frac{2A}{n\delta}}\qquad$
Under the same setting of Theorem 3.1, if $\varepsilon_{n}<\beta/4$ (equivalently $n>\frac{32AB^{2}}{\beta^{2}\delta}$), then for $\delta\in(0,1)$, with probability at least $1-\delta$ over $\vartheta_{n}$, the following (denoted by event $V$) hold simultaneously:
  1. 1.
    For $s$ with $|s|<\beta-\varepsilon_{n}$, $f_{S\mid\vartheta_{n}}\left(s\mid\vartheta_{n}\right)\geq 2f_{\min}$.
  2. 2.
    $\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\leq\varepsilon_{n}<\beta/4$.
Proof.
By the definition of $S$ in (8),
$\mathbb{P}\left[S\leq s|X,\vartheta_{n}\right]=\mathbb{P}\left[\begin{array}[]{l}\left.t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-s\leq Y\leq t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+s]\right.\\ \text{ and }s\geq\frac{t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)}{2}\end{array}\middle|X,\vartheta_{n}\right]$
Hence,
$F_{S|X,\vartheta_{n}}\left(s\right)=\begin{cases}0,\;&\text{if }s<\frac{t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)}{2},\\ F_{Y|X,\vartheta_{n}}\left(t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+s\right)\\ \qquad-F_{Y|X,\vartheta_{n}}\left(t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-s\right),\;&\text{otherwise.}\end{cases}$ (39)
We now show that with high probability, it holds for $s$ in the neighbourhood of 0 that
$s\geq\frac{t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)}{2},\quad t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+s\in\mathcal{Y},\quad t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-s\in\mathcal{Y}$
Let $y_{\max}:=\sup\{y\in\mathcal{Y}\}$ and $y_{\min}:=\inf\{y\in\mathcal{Y}\}$. Then, under Assumption 3.3, $y_{\max}>y_{\min}$.
$q_{\alpha/2}\left(Y\mid X=x\right),q_{1-\alpha/2}\left(Y\mid X=x\right)\in[y_{\min},y_{\max}],$
$q_{\alpha/2}\left(Y\mid X=x\right)-y_{\min}\geq\frac{\alpha}{2f_{\max}}\geq\beta,\qquad y_{\max}-q_{1-\alpha/2}\left(Y\mid X=x\right)\geq\frac{\alpha}{2f_{\max}}\geq\beta$
$\frac{q_{1-\alpha/2}\left(Y\mid X=x\right)-q_{\alpha/2}\left(Y\mid X=x\right)}{2}\geq\frac{1-\alpha}{2f_{\max}}\geq\beta$
By Theorem 3.1, $\mathbb{E}_{\theta_{n}}\left[\|\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\|_{2}^{2}\right]\leq\frac{A}{n}$ for $\gamma\in\{\frac{\alpha}{2},1-\frac{\alpha}{2}\}$. By Markov’s inequality,
$\mathbb{P}\left[\|\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\|_{2}\leq\sqrt{\frac{2A}{n\delta}}\right]\geq 1-\frac{\delta}{2}$
Applying the union bound, we have
$\mathbb{P}\left[\max_{\gamma\in\{\frac{\alpha}{2},1-\frac{\alpha}{2}\}}\|\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\|_{2}\leq\sqrt{\frac{2A}{n\delta}}\right]\geq 1-\delta$
Since for each $x\in\mathcal{X}$,
$\mathcal{E}_{\gamma}\left(x,\theta_{n}\left(\gamma\right)\right)$ $=\left|t_{\gamma}\left(x;\theta_{n}\left(\gamma\right)\right)-t_{\gamma}\left(x;\theta^{*}\left(\gamma\right)\right)\right|=\left|\left(\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\right)^{\top}x\right|$
$\leq\left\|\left(\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\right)\right\|_{2}\left\|x\right\|_{2}\leq B\left\|\left(\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\right)\right\|_{2}$
we have that with probability at least $1-\delta$,
$\sup_{x}\ \Delta\left(x,\vartheta_{n}\right)$ $\leq B\max_{\gamma\in\{\frac{\alpha}{2},1-\frac{\alpha}{2}\}}\|\theta_{n}\left(\gamma\right)-\theta^{*}\left(\gamma\right)\|_{2}\leq B\sqrt{\frac{2A}{n\delta}}=:\varepsilon_{n}$
and by Proposition B.5, it also holds that
$\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\leq\varepsilon_{n}$ (40)
Then, w.p. $\geq 1-\delta$, for any $x\in\mathcal{X}$,
$t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)$ $\geq q_{\alpha/2}\left(Y\mid X=x\right)-\Delta\left(x,\vartheta_{n}\right)\geq y_{\min}+\beta-\varepsilon_{n}$
$t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)$ $\leq q_{1-\alpha/2}\left(Y\mid X=x\right)+\Delta\left(x,\vartheta_{n}\right)\leq y_{\max}-\beta+\varepsilon_{n}$
$\frac{t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)}{2}$ $\geq\frac{q_{1-\alpha/2}\left(Y\mid X=x\right)-q_{\alpha/2}\left(Y\mid X=x\right)}{2}-\Delta\left(x,\vartheta_{n}\right)\geq\beta-\varepsilon_{n}$
The last inequality above shows that with high probability, quantile crossing will not occur given $n$ is large enough.
In this case, for any $s$ with $|s|<r_{n}:=\beta-\varepsilon_{n}$, we have $\forall x\in\mathcal{X}$,
$t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-s>y_{\min}+\beta-\varepsilon_{n}-r_{n}\geq y_{\min}$
$t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-s<q_{\alpha/2}\left(Y\mid X=x\right)+\varepsilon_{n}+r_{n}\leq q_{1-\alpha/2}\left(Y\mid X=x\right)+\beta\leq y_{\max}$
$t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+s<y_{\max}-\beta+\varepsilon_{n}+r_{n}\leq y_{\max}$
$t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+s>q_{1-\alpha/2}\left(Y\mid X=x\right)-\varepsilon_{n}-r_{n}\geq q_{\alpha/2}\left(Y\mid X=x\right)-\beta\geq y_{\min}$
$s\geq-|s|\geq-r_{n}=\varepsilon_{n}-\beta\geq\frac{t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)}{2}$
Since $\mathcal{Y}$ is an interval,
$t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-s\in\mathcal{Y},\qquad t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+s\in\mathcal{Y}$
Therefore, by (39), conditioning on $\vartheta_{n}$, for $s$ with $|s|<r_{n}=\beta-\varepsilon_{n}$,
$f_{S\mid\vartheta_{n}}\left(s\mid\vartheta_{n}\right)$ $=\mathbb{E}_{X|\vartheta_{n}}\left[\right$. $\left.f_{Y|X,\vartheta_{n}}\left(t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-s\mid X,\vartheta_{n}\right)\right$.
$\left.+f_{Y|X,\vartheta_{n}}\left(t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+s\mid X,\vartheta_{n}\right)\right]$
$\geq 2f_{\min}$
Suppose $n>\frac{32AB^{2}}{\beta^{2}\delta}$, which is equivalent to $\varepsilon_{n}<\beta/4$. Then, $r_{n}=\beta-\varepsilon_{n}\geq 3\beta/4\geq\varepsilon_{n}$. By (40), $\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\leq\beta-\varepsilon_{n}$. ∎
The proof of Proposition B.7 also relies on the following useful lemma.
Lemma B.9.
Let $F$ be a c.d.f. with p.d.f. $f$. Suppose there exists an interval $\mathcal{I}\in\mathbb{R}$ and a constant $c_{0}>0$ such that $f(s)\geq c_{0}$ for all $s\in\mathcal{I}$. For $p\in\left(0,1\right)$, $q_{p}:=\inf\{u:F\left(u\right)\geq p\}\in\mathcal{I}$, define $r_{0}:=\min\{q_{p}-\inf\mathcal{I},\sup\mathcal{I}-q_{p}\}\geq 0$. Then, for any $p^{\prime}$ such that $|p^{\prime}-p|<c_{0}r_{0}$, it holds that $q_{p^{\prime}}\in\mathcal{I}$, and $|q_{p^{\prime}}-q_{p}|\leq\frac{|p^{\prime}-p|}{c_{0}}$.
Proof.
By assumption,
$F\left(q_{p}-r_{0}\right)\leq F\left(q_{p}\right)-c_{0}r_{0}=p-c_{0}r_{0}$
$F\left(q_{p}+r_{0}\right)\geq F\left(q_{p}\right)+c_{0}r_{0}=p+c_{0}r_{0}$
Since $|p^{\prime}-p|<c_{0}r_{0}$, either $p\leq p^{\prime}<p+c_{0}r_{0}$ or $p^{\prime}\leq p<p^{\prime}+c_{0}r_{0}$. If $p\leq p^{\prime}<p+c_{0}r_{0}$, then $p\leq p^{\prime}<F\left(q_{p}+r_{0}\right)$. Since $F$ is non-decreasing, $q_{p}\leq q_{p^{\prime}}<q_{p}+r_{0}$. Similarly, if $p-c_{0}r_{0}<p^{\prime}\leq p$, then $F\left(q_{p}-r_{0}\right)<p^{\prime}\leq p$, and $q_{p}-r_{0}<q_{p^{\prime}}\leq q_{p}$. Hence, $q_{p^{\prime}}\in\mathcal{I}$, and $|q_{p^{\prime}}-q_{p}|\leq\frac{|p^{\prime}-p|}{c_{0}}$. ∎
With Proposition B.8, we apply Lemma B.9 and get Lemma B.10.
Lemma B.10.
Under the same setting of Proposition B.8, if the event in Proposition B.8 occurs, and if $m>\frac{4}{f_{\min}\beta}$, then it holds that
  • $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\leq\beta/2$;
  • $f_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)\right)\geq 2f_{\min}$;
  • $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|\leq\frac{1}{f_{\min}m}$.
Proof.
For simplicity, in the proof we denote $q_{p}\left(S\mid\vartheta_{n}\right)$ by $q_{p}$.
$\left(1-\alpha\right)\left(m+1\right)\leq\lceil\left(1-\alpha\right)\left(m+1\right)\rceil<\left(1-\alpha\right)\left(m+1\right)+1$
$\Rightarrow\;$ $\left(1-\alpha\right)\left(m+1\right)-\left(1-\alpha\right)m\leq\lceil\left(1-\alpha\right)\left(m+1\right)\rceil-\left(1-\alpha\right)m<\left(1-\alpha\right)\left(m+1\right)+1-\left(1-\alpha\right)m$
$\Rightarrow\;$ $0<\frac{1-\alpha}{m}\leq|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|<\frac{2-\alpha}{m}<\frac{2}{m}$
Since $\varepsilon_{n}<\beta/4$, from Proposition B.8, with probability at least $1-\delta$, for $s$ with $|s|<3\beta/4$, $f_{S\mid\vartheta_{n}}\left(s\mid\vartheta_{n}\right)\geq 2f_{\min}$, and $|q_{1-\alpha}|<\beta/4$. In this case, $r_{0}:=\min\{q_{1-\alpha}+3\beta/4,3\beta/4-q_{1-\alpha}\}>\beta/2$. If $m>\frac{4}{f_{\min}\beta}$, then $|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|<\frac{2}{m}<2f_{\min}\frac{\beta}{4}<2f_{\min}\frac{\beta}{2}<2f_{\min}\cdot r_{0}$. Then by Lemma B.9, $|q_{\left(1-\alpha\right)_{m}}|\leq 3\beta/4$, $f_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)\right)\geq 2f_{\min}$, and $|q_{\left(1-\alpha\right)_{m}}-q_{1-\alpha}|<\frac{|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|}{2f_{\min}}<\frac{1}{f_{\min}m}\leq\beta/4$. Hence, $|q_{\left(1-\alpha\right)_{m}}|\leq|q_{1-\alpha}|+|q_{\left(1-\alpha\right)_{m}}-q_{1-\alpha}|<\beta/4+\beta/4=\beta/2$. ∎
Notice that $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|$ is bounded by $2R$. Let $V$ denote the event in Proposition B.8, and $V^{c}$ its complement. Then, by Lemma B.10,
$\mathbb{E}_{\vartheta_{n}}\left[|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|\right]$
$=\mathbb{P}[V]\cdot\mathbb{E}_{\vartheta_{n}}\left[|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|\mid V\right]$
$\quad+\mathbb{P}\left[V^{c}\right]\cdot\mathbb{E}_{\vartheta_{n}}\left[|q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)-q_{1-\alpha}\left(S\mid\vartheta_{n}\right)|\mid V^{c}\right]$
$\leq\frac{1}{f_{\min}m}+2R\delta$
Picking $\delta=\frac{33AB^{2}}{\beta^{2}n}$ completes the proof of Proposition B.7.
B.4 Proof of Proposition B.11
Proposition B.11 (Empirical score-quantile concentration).
In CQR, Suppose Assumptions 3.13.3 hold, if $m>8H/\min\{\alpha,1-\alpha\}$ for $H$ in (12), then
$\mathbb{E}_{\vartheta_{n},\mathcal{D}_{\text{cal}}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\right]$
$\quad\leq\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2m}}+4R\exp{\left(-\frac{\min\{\alpha^{2},(1-\alpha)^{2}\}f_{\min}^{2}}{8f_{\max}^{2}}m\right)}+\frac{1056Rf_{\max}^{3}\lambda_{\max}^{2}B^{2}d}{\min\{\alpha^{2},(1-\alpha)^{2}\}\lambda_{\min}^{4}f_{\min}^{2}n}.$
To prove Proposition B.11, we first prove the following lemma:
Lemma B.12.
Under the same setting of Lemma B.10, if the high probability event $V$ in Proposition B.8 occurs, for any $u\in[0,\beta/4]$, if
$\sup_{s}\left|F_{S\mid\vartheta_{n}}\left(s\right)-\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(s\right)\right|\leq 2f_{\min}u$
then $|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\leq u$.
Proof.
For simplicity, in the proof we denote $q_{p}\left(S\mid\vartheta_{n}\right)$ by $q_{p}$. By Lemma B.10, for $u\in[0,\beta/4]$, $|q_{\left(1-\alpha\right)_{m}}-u|\leq 3\beta/4$ and $|q_{\left(1-\alpha\right)_{m}}+u|\leq 3\beta/4$. Hence, in this case,
$F_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)\leq F_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}\right)-2f_{\min}u=\left(1-\alpha\right)_{m}-2f_{\min}u$
$F_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)\geq F_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}\right)+2f_{\min}u=\left(1-\alpha\right)_{m}+2f_{\min}u$
By assumption,
$\left|F_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)-\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)\right|\leq 2f_{\min}u$
$\left|F_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)-\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)\right|\leq 2f_{\min}u$
Then
$\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)\leq\left(1-\alpha\right)_{m},\qquad\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)\geq\left(1-\alpha\right)_{m}$
Since $\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}$ is non-decreasing, we have
$\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right):=\inf\{u^{\prime}\in\mathcal{S}_{m}:\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(u^{\prime}\right)\geq\left(1-\alpha\right)_{m}\}\in\left[q_{\left(1-\alpha\right)_{m}}-u,q_{\left(1-\alpha\right)_{m}}+u\right]$
where $\mathcal{S}_{m}$ is the set of scores of the calibration data.
Then, $|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\leq u$. ∎
Lemma B.13 (Dvoretzky–Kiefer–Wolfowitz Inequality (Dvoretzky et al., 1956; Massart, 1990)).
Given a natural number $m$, let $X_{1},\dots,X_{m}$ be real-valued i.i.d. random variables with c.d.f. $F\left(\cdot\right)$. Let $F^{\left(m\right)}$ denote the associated empirical distribution function defined by
$F^{\left(m\right)}\left(x\right)=\frac{1}{m}\sum_{j=1}^{m}\mathbbm{1}\{X_{j}\leq x\},\qquad x\in\mathbb{R}$
Then,
$\mathbb{P}\left[\sup_{x\in\mathbb{R}}\left|F^{\left(m\right)}\left(x\right)-F\left(x\right)\right|>\varepsilon\right]$ $\leq 2e^{-2m\varepsilon^{2}}$ $\forall\varepsilon\geq 0$
By the Dvoretzky–Kiefer–Wolfowitz Inequality (Lemma B.13),
$\mathbb{P}\left[\sup_{s}\left|F_{S\mid\vartheta_{n}}\left(s\right)-\hat{F}^{\left(m\right)}_{S\mid\vartheta_{n}}\left(s\right)\right|\geq 2f_{\min}u\right]\leq 2\exp{\left(-8mf_{\min}^{2}u^{2}\right)}$
Thus, by Lemma B.12, given that the event $V$ occurs,
$\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\geq u\;\Big|\;V\right]\leq 2\exp{\left(-8mf_{\min}^{2}u^{2}\right)},\quad u\in[0,\beta/4].$
Specifically,
$\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\geq\beta/4\;\Big|\;V\right]\leq 2\exp{\left(-8mf_{\min}^{2}(\beta/4)^{2}\right)}$
Then, for any $u>\beta/4$,
$\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\geq u\;\Big|\;V\right]\leq 2\exp{\left(-8mf_{\min}^{2}(\beta/4)^{2}\right)}$
Since $|S|\leq R$, $|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\leq 2R$. By the layer cake representation of the expectation of a non-negative random variable $Z$, which is $\mathbb{E}[Z]=\int_{0}^{\infty}\mathbb{P}[Z\geq u]\ du$,
$\mathbb{E}_{\vartheta_{n},\mathcal{D}_{\mathrm{cal}}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\;\Big|\;V\right]$
$=\int_{0}^{2R}\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\geq u\;\Big|\;V\right]du$
$\leq\int_{0}^{\beta/4}2\exp{\left(-8mf_{\min}^{2}u^{2}\right)}du+\int_{\beta/4}^{2R}2\exp{\left(-8mf_{\min}^{2}(\beta/4)^{2}\right)}du$
$\leq 2\int_{0}^{\infty}\exp{\left(-8mf_{\min}^{2}u^{2}\right)}du+4R\exp{\left(-8f_{\min}^{2}(\beta/4)^{2}m\right)}$
$=\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2m}}+4R\exp{\left(-\frac{1}{2}f_{\min}^{2}\beta^{2}m\right)}$
Therefore, we have
$\mathbb{E}_{\vartheta_{n},\mathcal{D}_{\mathrm{cal}}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\right]$
$\leq\mathbb{P}\left[V\right]\cdot\mathbb{E}_{\vartheta_{n}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\vartheta_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\vartheta_{n}\right)|\;\Big|\;V\right]+\mathbb{P}\left[V^{c}\right]\cdot 2R$
$\leq\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2m}}+4R\exp{\left(-\frac{1}{2}f_{\min}^{2}\beta^{2}m\right)}+2R\delta$
Picking $\delta=\frac{33AB^{2}}{\beta^{2}n}$ completes the proof of Proposition B.11.
B.5 Proof of Theorem 3.2
See 3.2
Proof.
By the definition of the prediction set (9),
$\left|\mathcal{C}(x)\right|$ $=\max\left\{0,\;t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)+2\hat{q}_{(1-\alpha)_{m}}\right\}$
$\leq\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)+2\hat{q}_{(1-\alpha)_{m}}\right|$
We further bound the right hand side by
$\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)+2\hat{q}_{(1-\alpha)_{m}}\right|$
$=\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)+t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)+t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right$.
$\qquad\left.+2\hat{q}_{(1-\alpha)_{m}}\right|$
$\leq\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)\right|+\left|t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right|+2\left|\hat{q}_{(1-\alpha)_{m}}\right|$
$\qquad+\left|t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right|$
$=\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)\right|+\left|t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right|+2\left|\hat{q}_{(1-\alpha)_{m}}\right|$
$\qquad+\left(t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right),$
where the last equality follows because
$t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)=q_{1-\alpha/2}\left(Y\mid X\right)\geq q_{\alpha/2}\left(Y\mid X\right)=t_{\alpha/2}\left(x;\underline{\theta}^{*}\right).$
Hence,
$\left|\mathcal{C}(X)\right|-\left(t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right)$
$\qquad\leq\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)\right|+\left|t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right|+2\left|\hat{q}_{(1-\alpha)_{m}}\right|$
We also have
$-\left(\left|\mathcal{C}(X)\right|-\left(t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right)\right)$
$=\left(t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right)-\max\left\{0,\;t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)+2\hat{q}_{(1-\alpha)_{m}}\right\}$
$\leq t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)+t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-2\hat{q}_{(1-\alpha)_{m}}$
$\leq\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)\right|+\left|t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right|+2\left|\hat{q}_{(1-\alpha)_{m}}\right|$
Therefore,
$\left|\left|\mathcal{C}(X)\right|-\left(t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right)\right|$
$\qquad\leq\left|t_{1-\alpha/2}\left(x;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(x;\bar{\theta}^{*}\right)\right|+\left|t_{\alpha/2}\left(x;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(x;\underline{\theta}^{*}\right)\right|+2\left|\hat{q}_{(1-\alpha)_{m}}\right|$
Hence, for test sample $(X,Y)$,
$\mathbb{E}_{X,\vartheta_{n},\mathcal{D}_{\text{cal}}}\left[\left|\left|\mathcal{C}(X)\right|-t_{1-\alpha/2}\left(X;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(X;\underline{\theta}^{*}\right)\right|\right]$
$\leq\mathbb{E}_{X,\vartheta_{n}}\left[\left|t_{1-\alpha/2}\left(X;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(X;\bar{\theta}^{*}\right)\right|\right]+\mathbb{E}_{X,\vartheta_{n}}\left[\left|t_{\alpha/2}\left(X;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(X;\underline{\theta}^{*}\right)\right|\right]$
$\hskip 18.49988pt+2\mathbb{E}_{\vartheta_{n},\mathcal{D}_{\text{cal}}}\left[\left|\hat{q}_{(1-\alpha)_{m}}\left(S_{m}\mid\vartheta_{n}\right)\right|\right]$
$\leq\mathbb{E}_{X,\vartheta_{n}}\left[\left|t_{1-\alpha/2}\left(X;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(X;\bar{\theta}^{*}\right)\right|\right]+\mathbb{E}_{X,\vartheta_{n}}\left[\left|t_{\alpha/2}\left(X;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(X;\underline{\theta}^{*}\right)\right|\right]$
$\hskip 18.49988pt+2\mathbb{E}_{\vartheta_{n}}\left[\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\right]+2\mathbb{E}_{\vartheta_{n},\mathcal{D}_{\text{cal}}}\left[\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)-\hat{q}_{(1-\alpha)_{m}}\left(S_{m}\mid\vartheta_{n}\right)\right|\right]$
$\leq\mathbb{E}_{X,\vartheta_{n}}\left[\left|t_{1-\alpha/2}\left(X;\bar{\theta}_{n}\right)-t_{1-\alpha/2}\left(X;\bar{\theta}^{*}\right)\right|\right]+\mathbb{E}_{X,\vartheta_{n}}\left[\left|t_{\alpha/2}\left(X;\underline{\theta}_{n}\right)-t_{\alpha/2}\left(X;\underline{\theta}^{*}\right)\right|\right]$
$\hskip 18.49988pt+2\mathbb{E}_{\vartheta_{n}}\left[\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)\right|\right]+2\mathbb{E}_{\vartheta_{n}}\left[\left|q_{1-\alpha}\left(S\mid\vartheta_{n}\right)-q_{(1-\alpha)_{m}}\left(S\mid\vartheta_{n}\right)\right|\right]$
$\hskip 18.49988pt+2\mathbb{E}_{\vartheta_{n},\mathcal{D}_{\text{cal}}}\left[\left|q_{(1-\alpha)_{m}}\left(S\mid\vartheta_{n}\right)-\hat{q}_{(1-\alpha)_{m}}\left(S_{m}\mid\vartheta_{n}\right)\right|\right]$
By Theorem 3.1,
$\mathbb{E}_{X,\theta_{n}}\left[\left|t_{\gamma}\left(X;\theta_{n}\left(\gamma\right)\right)-t_{\gamma}\left(X;\theta^{*}\left(\gamma\right)\right)\right|\right]$ $\leq\sqrt{\mathbb{E}_{X,\theta_{n}}\left[\left(t_{\gamma}\left(X;\theta_{n}\left(\gamma\right)\right)-t_{\gamma}\left(X;\theta^{*}\left(\gamma\right)\right)\right)^{2}\right]}$
$\leq\frac{2\lambda_{\max}\sqrt{f_{\max}d}}{\lambda_{\min}f_{\min}\sqrt{\lambda_{\min}n}}$
By Proposition B.5,B.7,B.11,
$\mathbb{E}_{X,\vartheta_{n},\mathcal{D}_{\text{cal}}}\left[\left|\left|\mathcal{C}(X)\right|-t_{1-\alpha/2}\left(X;\bar{\theta}^{*}\right)-t_{\alpha/2}\left(X;\underline{\theta}^{*}\right)\right|\right]$
$\leq\left(\frac{4\lambda_{\max}\sqrt{f_{\max}d}}{\lambda_{\min}f_{\min}\sqrt{\lambda_{\min}}}+\frac{2B\ \lambda_{\max}\sqrt{2f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\right)\sqrt{\frac{1}{n}}+\frac{1}{f_{\min}m}$
$\qquad+\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2m}}+4R\exp{\left(-\frac{1}{2}f_{\min}^{2}\beta^{2}m\right)}+\frac{66AB^{2}R}{\beta^{2}n}$
$=\left(\frac{4\lambda_{\max}\sqrt{f_{\max}d}}{\lambda_{\min}f_{\min}\sqrt{\lambda_{\min}}}+\frac{2B\ \lambda_{\max}\sqrt{2f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\right)\sqrt{\frac{1}{n}}+\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2}}\sqrt{\frac{1}{m}}+\frac{1}{f_{\min}m}$
$\qquad+4R\exp{\left(-\frac{\min\{\alpha^{2},(1-\alpha)^{2}\}f_{\min}^{2}}{8f_{\max}^{2}}m\right)}+\frac{1056\lambda_{\max}^{2}f_{\max}^{3}B^{2}R}{\min\{\alpha^{2},(1-\alpha)^{2}\}\lambda_{\min}^{4}f_{\min}^{2}n}$ (41)
This completes the proof of Theorem 3.2. ∎
Appendix C Proofs of Results in CMR
To prove Theorem 4.1, the goal is to upper bound
$\mathbb{E}_{X,\check{\theta}_{n},\mathcal{D}_{\mathrm{cal}}}\left[\left|2\ \hat{q}_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1-\alpha/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|\right]$
$=2\ \mathbb{E}_{X,\check{\theta}_{n},\mathcal{D}_{\mathrm{cal}}}\left[\left|\hat{q}_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|\right]$
Further decompose it, and we have
$\left|\hat{q}_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|$
$=\left|\hat{q}_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)+q_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)\right$.
$\qquad\left.+q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|$
$\leq\left|\hat{q}_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)\right|+\left|q_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)\right|$
$\qquad+\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|$
Thus, the expectation is decomposed into three parts as follows, and we will upper bound each of them in Proposition C.4, C.3, and C.1:
$\mathbb{E}_{X,\check{\theta}_{n},\mathcal{D}_{\mathrm{cal}}}\left[\left|2\ \hat{q}_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1-\alpha/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|\right]$
$=2\ \mathbb{E}_{\check{\theta}_{n},\mathcal{D}_{\mathrm{cal}}}\left[\left|\hat{q}_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)\right|\right]$
$\qquad+2\ \mathbb{E}_{\check{\theta}_{n}}\left[\left|q_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)\right|\right]$
$\qquad+2\ \mathbb{E}_{X,\check{\theta}_{n}}\left[\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|\right]$
$\leq\frac{\sqrt{\pi}}{f_{\min}\sqrt{2m}}+8R\exp{\left(-\frac{f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}}{8f_{\max}^{2}}m\right)}+\frac{2056R\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}n}$
$\qquad+\frac{2}{f_{\min}m}+\frac{4B\ \lambda_{\max}\sqrt{f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\sqrt{\frac{1}{n}}$ (42)
To proceed, we define some random variables for simplicity.
$\Delta\left(X,\check{\theta}_{n}\right)$ $:=\left|t_{1/2}\left(X;\check{\theta}_{n}\right)-t_{1/2}\left(X;\check{\theta}^{*}\right)\right|\geq 0$ (43)
$S^{*}\left(X,Y\right)$ $:=\left|q_{1/2}(Y\mid X)-Y\right|$ (44)
$M\left(\check{\theta}_{n}\right)$ $:=\left\|\left(\check{\theta}_{n}-\check{\theta}^{*}\right)\right\|_{2}$ (45)
C.1 Proof of Proposition C.1
Proposition C.1.
In CMR, suppose Assumption 4.2 holds, we have
$\left|q_{1-\alpha}\left(S\mid X,\check{\theta}_{n}\right)-\zeta\right|\leq B\cdot M\left(\check{\theta}_{n}\right)$ (46)
If Assumptions 4.1,3.2,3.3 further hold, then
$\mathbb{E}_{X,\check{\theta}_{n}}\left[\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|\right]\leq\frac{2B\ \lambda_{\max}\sqrt{f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\sqrt{\frac{1}{n}}$ (47)
Proof.
Notice that
$S\left(X,Y;\check{\theta}_{n}\right)$ $:=\left|t_{1/2}\left(X;\check{\theta}_{n}\right)-Y\right|$
$\leq\left|q_{1/2}(Y\mid X)-Y\right|+\left|t_{1/2}\left(X;\check{\theta}_{n}\right)-q_{1/2}(Y\mid X)\right|$
$=S^{*}\left(X,Y\right)+\Delta\left(X,\check{\theta}_{n}\right)$
Similarly, $S\left(X,Y;\check{\theta}_{n}\right)\geq S^{*}\left(X,Y\right)-\Delta\left(X,\check{\theta}_{n}\right)$. Hence,
$\left|S\left(X,Y;\check{\theta}_{n}\right)-S^{*}\left(X,Y\right)\right|\leq\Delta\left(X,\check{\theta}_{n}\right)\leq\|X\|_{2}\left\|\left(\check{\theta}_{n}-\check{\theta}^{*}\right)\right\|_{2}\leq B\cdot\left\|\left(\check{\theta}_{n}-\check{\theta}^{*}\right)\right\|_{2}$
Now we show that $q_{1-\alpha}\left(S^{*}\mid X\right)=q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)$. Note that given $X$,
$S^{*}\left(X,Y\right)\leq q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)$
$\Longleftrightarrow\;$ $-\left(q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)\right)\leq Y-q_{1/2}(Y\mid X)\leq q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)$
$\Longleftrightarrow\;$ $q_{\alpha/2}(Y\mid X)\leq Y\leq q_{1-\alpha/2}(Y\mid X)$
where the last step is from Assumption 4.2. Since $F_{Y\mid X}$ is continuous,
$\mathbb{P}\left[q_{\alpha/2}\left(Y\mid X\right)\leq Y\leq q_{1-\alpha/2}\left(Y\mid X\right)\mid X\right]=1-\alpha.$
Hence,
$\mathbb{P}[S^{*}\left(X,Y\right)\leq q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)|X]=1-\alpha.$
Let $q_{1-\alpha}\left(S^{*}\mid X\right)$ be the $\left(1-\alpha\right)$-quantile of $S^{*}$ given $X$. Since $X$ is given, and $F_{Y\mid X}$ is continuous, $F_{S^{*}|X}$ is continuous. Then, $q_{1-\alpha}\left(S^{*}\mid X\right)=q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)$.
Conditioned on $X,\check{\theta}_{n}$, $\Delta\left(X,\check{\theta}_{n}\right)$ is deterministic. Thus,
$\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\leq u\mid X,\check{\theta}_{n}\right]\geq\mathbb{P}\left[S^{*}\left(X,Y\right)+\Delta\left(X,\check{\theta}_{n}\right)\leq u\mid X,\check{\theta}_{n}\right]$
$\Rightarrow\;$ $\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\leq\Delta\left(X,\check{\theta}_{n}\right)+q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)\mid X,\check{\theta}_{n}\right]$
$\quad\geq\mathbb{P}\left[S^{*}\left(X,Y\right)\leq q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)\mid X\right]=1-\alpha$
Then, $q_{1-\alpha}\left(S\mid X,\check{\theta}_{n}\right)\leq\Delta\left(X,\check{\theta}_{n}\right)+q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)$. Similarly, we have
$\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\leq u\mid X,\check{\theta}_{n}\right]\leq\mathbb{P}\left[S^{*}\left(X,Y\right)-\Delta\left(X,\check{\theta}_{n}\right)\leq u\mid X,\check{\theta}_{n}\right]$
$\Rightarrow\;$ $\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\leq-\Delta\left(X,\check{\theta}_{n}\right)+q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)\mid X,\check{\theta}_{n}\right]$
$\quad\leq\mathbb{P}\left[S^{*}\left(X,Y\right)\leq q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)\mid X\right]=1-\alpha$
Then, $q_{1-\alpha}\left(S\mid X,\check{\theta}_{n}\right)\geq-\Delta\left(X,\check{\theta}_{n}\right)+q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)$. Thus, by Assumption 4.2,
$\left|q_{1-\alpha}\left(S\mid X,\check{\theta}_{n}\right)-\left(q_{1/2}(Y\mid X)-q_{\alpha/2}(Y\mid X)\right)\right|\leq\Delta\left(X,\check{\theta}_{n}\right)$
$\Longrightarrow\left|q_{1-\alpha}\left(S\mid X,\check{\theta}_{n}\right)-\zeta\right|\leq B\cdot M\left(\check{\theta}_{n}\right)$
Then we can remove the conditioning on $X$,
$\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\leq\zeta+B\cdot M\left(\check{\theta}_{n}\right)\mid\check{\theta}_{n}\right]$
$=\mathbb{E}_{X,Y|\check{\theta}_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\check{\theta}_{n}\right)\leq\zeta+B\cdot M\left(\check{\theta}_{n}\right)\right\}\mid\check{\theta}_{n}\right]$
$=\mathbb{E}_{X|\check{\theta}_{n}}\left[\mathbb{E}_{Y|X,\check{\theta}_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\check{\theta}_{n}\right)\leq\zeta+B\cdot M\left(\check{\theta}_{n}\right)\right\}\mid X,\check{\theta}_{n}\right]\mid\check{\theta}_{n}\right]$
$=\mathbb{E}_{X|\check{\theta}_{n}}\left[\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\leq\zeta+B\cdot M\left(\check{\theta}_{n}\right)\mid X,\check{\theta}_{n}\right]\mid\check{\theta}_{n}\right]$
$\geq\mathbb{E}_{X|\check{\theta}_{n}}\left[1-\alpha\mid\check{\theta}_{n}\right]=1-\alpha$
Hence, $q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)\leq\zeta+B\cdot M\left(\check{\theta}_{n}\right)$. And by similar arguments as below, $q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)\geq\zeta-B\cdot M\left(\check{\theta}_{n}\right)$. Specifically,
$\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\geq\zeta-B\cdot M\left(\check{\theta}_{n}\right)\mid\check{\theta}_{n}\right]$
$=\mathbb{E}_{X,Y|\check{\theta}_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\check{\theta}_{n}\right)\geq\zeta-B\cdot M\left(\check{\theta}_{n}\right)\right\}\mid\check{\theta}_{n}\right]$
$=\mathbb{E}_{X|\check{\theta}_{n}}\left[\mathbb{E}_{Y|X,\check{\theta}_{n}}\left[\mathbbm{1}\left\{S\left(X,Y;\check{\theta}_{n}\right)\geq\zeta-B\cdot M\left(\check{\theta}_{n}\right)\right\}\mid X,\check{\theta}_{n}\right]\mid\check{\theta}_{n}\right]$
$=\mathbb{E}_{X|\check{\theta}_{n}}\left[\mathbb{P}\left[S\left(X,Y;\check{\theta}_{n}\right)\geq\zeta-B\cdot M\left(\check{\theta}_{n}\right)\mid X,\check{\theta}_{n}\right]\mid\check{\theta}_{n}\right]$
$\leq\mathbb{E}_{X|\check{\theta}_{n}}\left[1-\alpha\mid\check{\theta}_{n}\right]=1-\alpha$
Therefore, $\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\leq B\cdot M\left(\check{\theta}_{n}\right)$.
Then, by Theorem 3.1,
$\mathbb{E}_{\check{\theta}_{n}}\left[\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\right]$ $\leq B\cdot\mathbb{E}_{\check{\theta}_{n}}\left[M\left(\check{\theta}_{n}\right)\right]\leq B\ \sqrt{\mathbb{E}_{\check{\theta}_{n}}\left[\left\|\left(\underline{\theta}_{n}-\underline{\theta}^{*}\right)\right\|_{2}^{2}\right]}$
$\leq B\ \sqrt{\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{4}f_{\min}^{2}n}}=\frac{2B\ \lambda_{\max}\sqrt{f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\sqrt{\frac{1}{n}}$
i.e.,
$\mathbb{E}_{X,\check{\theta}_{n}}\left[\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\left(q_{1/2}\left(Y\mid X\right)-q_{\alpha/2}\left(Y\mid X\right)\right)\right|\right]\leq\frac{2B\ \lambda_{\max}\sqrt{f_{\max}d}}{\lambda_{\min}^{2}f_{\min}}\sqrt{\frac{1}{n}}$
C.2 Proof of Proposition C.2
Proposition C.2.
In CMR, suppose Assumption 4.1,3.2,3.3,4.2 hold. Define
$\beta:=\min\left\{\frac{\alpha}{2f_{\max}},\ \frac{1-\alpha}{2f_{\max}}\right\}\qquad\varepsilon_{n}:=B\sqrt{\frac{A}{n\delta}}$ (48)
If $\varepsilon_{n}<\beta/4$, then with probability at least $1-\delta$, for any $s$ such that for $s\in\mathcal{I}:=\{s\in\mathbb{R}:|s-\zeta|\leq\beta-\varepsilon_{n}\}$,
$f_{S|\check{\theta}_{n}}\left(s\right)\geq 2f_{\min}$
and $\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\leq\varepsilon_{n}\leq\beta-\varepsilon_{n}$.
Proof.
By the definition of $S$,
$\mathbb{P}\left[S\leq s|X,\check{\theta}_{n}\right]=\mathbb{P}\left[t_{1/2}\left(X;\check{\theta}_{n}\right)-s\leq Y\leq t_{1/2}\left(X;\check{\theta}_{n}\right)+s\mid X,\check{\theta}_{n}\right]$
Hence,
$F_{S|X,\check{\theta}_{n}}\left(s\right)=F_{Y|X,\check{\theta}_{n}}\left(t_{1/2}\left(x;\check{\theta}_{n}\right)+s\right)-F_{Y|X,\check{\theta}_{n}}\left(t_{1/2}\left(x;\check{\theta}_{n}\right)-s\right)$ (49)
We now show that with high probability, it holds for $s$ in the neighbourhood of $\zeta$ that
$t_{1/2}\left(x;\check{\theta}_{n}\right)+s\in\mathcal{Y},\quad t_{1/2}\left(x;\check{\theta}_{n}\right)-s\in\mathcal{Y}$
By Theorem 3.1, $\mathbb{E}_{\check{\theta}_{n}}\left[\|\check{\theta}_{n}-\check{\theta}^{*}\|_{2}^{2}\right]\leq\frac{A}{n}$ for $A:=\frac{4\lambda_{\max}^{2}f_{\max}d}{\lambda_{\min}^{4}f_{\min}^{2}}$. By Markov’s inequality,
$\mathbb{P}\left[\|\check{\theta}_{n}-\check{\theta}^{*}\|_{2}\leq\sqrt{\frac{A}{n\delta}}\right]\geq 1-\delta$
Hence, with probability at least $1-\delta$,
$\sup_{x}\ \Delta\left(x,\check{\theta}_{n}\right)$ $\leq B\|\check{\theta}_{n}-\check{\theta}^{*}\|_{2}\leq B\sqrt{\frac{A}{n\delta}}=:\varepsilon_{n}$
In this case, by (46),
$\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\leq\varepsilon_{n}$ (50)
Then, for every $s$ such that $|s-\zeta|\leq\beta-\varepsilon_{n}$, i.e., $s\in\mathcal{I}$, it holds that
$t_{1/2}\left(x;\check{\theta}_{n}\right)+s\leq q_{1/2}\left(Y|X\right)+\varepsilon_{n}+\zeta+\beta-\varepsilon_{n}=q_{1/2}\left(Y|X\right)+\zeta+\beta=q_{1-\alpha/2}(Y|X)+\beta\leq y_{\max}$
$t_{1/2}\left(x;\check{\theta}_{n}\right)+s\geq q_{1/2}\left(Y|X\right)-\varepsilon_{n}+\zeta-\beta+\varepsilon_{n}=q_{1/2}\left(Y|X\right)+\zeta-\beta=q_{1-\alpha/2}\left(Y|X\right)-\beta\geq y_{\min}$
$t_{1/2}\left(x;\check{\theta}_{n}\right)-s\leq q_{1/2}\left(Y|X\right)+\varepsilon_{n}-\zeta+\beta-\varepsilon_{n}=q_{1/2}\left(Y|X\right)-\zeta+\beta=q_{\alpha/2}(Y|X)+\beta\leq y_{\max}$
$t_{1/2}\left(x;\check{\theta}_{n}\right)-s\geq q_{1/2}\left(Y|X\right)-\varepsilon_{n}-\zeta-\beta+\varepsilon_{n}=q_{1/2}\left(Y|X\right)-\zeta-\beta=q_{\alpha/2}\left(Y|X\right)-\beta\geq y_{\min}$
Thus, $t_{1/2}\left(x;\check{\theta}_{n}\right)+s\in\mathcal{Y},t_{1/2}\left(x;\check{\theta}_{n}\right)-s\in\mathcal{Y}$.
By (49), if $\varepsilon_{n}<\beta/4$, then with probability at least $1-\delta$, we have for any $s$ such that $|s-\zeta|\leq\beta-\varepsilon_{n}$,
$f_{S|X,\check{\theta}_{n}}\left(s\right)=f_{Y|X,\check{\theta}_{n}}\left(t_{1/2}\left(x;\check{\theta}_{n}\right)+s\right)+f_{Y|X,\check{\theta}_{n}}\left(t_{1/2}\left(x;\check{\theta}_{n}\right)-s\right)\geq 2f_{\min}$ (51)
Since $\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\leq\varepsilon_{n}\leq\beta-\varepsilon_{n}<\frac{3}{4}\beta$, after taking expectation over $X$, we have $f_{S|\check{\theta}_{n}}\left(q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right)\geq 2f_{\min}$. ∎
C.3 Proof of Proposition C.3
Proposition C.3.
In CMR, suppose Assumption 4.1,3.2,3.3,4.2 hold. If
$m>\frac{8f_{\max}}{f_{\min}\min\{\alpha,(1-\alpha)\}}.$ (52)
then
$\mathbb{E}_{\check{\theta}_{n}}\left[\left|q_{(1-\alpha)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)\right|\right]\leq\frac{1}{f_{\min}m}+\frac{514R\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}n}$ (53)
and if furthermore $n>\frac{256\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}\delta}$, then with probability at least $1-\delta$,
$|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)|\leq\frac{1}{f_{\min}m}<\frac{\beta}{4}.$
Proof.
Notice that
$0<\frac{1-\alpha}{m}\leq|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|<\frac{2-\alpha}{m}<\frac{2}{m}$
If let $m>\frac{4}{\beta f_{\min}}$ for $\beta$ defined in (48), then
$|\left(1-\alpha\right)_{m}-\left(1-\alpha\right)|<\frac{2}{m}<2f_{\min}\cdot\frac{\beta}{4}$
According to Lemma B.9, since $\left|q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)-\zeta\right|\leq\varepsilon_{n}<\frac{\beta}{4}$ by Proposition C.2, the distance from $\mathcal{I}^{c}$ is $r_{0}>\frac{\beta}{2}$. Thus, by Lemma B.9, $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)|\leq\frac{1}{f_{\min}m}<\frac{\beta}{4}$, and hence, $|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-\zeta|<\frac{\beta}{2}$.
Therefore, if $\varepsilon_{n}<\beta/4$ and $m>\frac{4}{f_{\min}\beta}$, then
$\mathbb{E}_{\check{\theta}_{n}}\left[|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)|\right]\leq\frac{1}{f_{\min}m}+2R\delta$
Taking $\delta=\frac{257\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}n}$, and we get
$\mathbb{E}_{\check{\theta}_{n}}\left[|q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)-q_{1-\alpha}\left(S\mid\check{\theta}_{n}\right)|\right]\leq\frac{1}{f_{\min}m}+\frac{514R\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}n}$
C.4 Proof of Proposition C.4
Proposition C.4.
In CMR, suppose Assumption 4.1,3.2,3.3,4.2 hold. If
$m>\frac{8H}{\min\{\alpha,(1-\alpha)\}}.$ (54)
for $H$ in (12), then
$\mathbb{E}_{\check{\theta}_{n},\mathcal{D}_{\mathrm{cal}}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\right]$
$\qquad\leq\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2m}}+4R\exp{\left(-\frac{f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}}{8f_{\max}^{2}}m\right)}+\frac{514Rf_{\max}^{3}\lambda_{\max}^{2}B^{2}d}{\min\{\alpha^{2},(1-\alpha)^{2}\}\lambda_{\min}^{4}f_{\min}^{2}n}.$
The proof of Proposition C.4 is essentially the same as the proof of Proposition B.11. We include here for completeness.
Proof.
Lemma C.5.
In CMR, under the same setting of Proposition C.2, if the high probability event $V$ in Proposition C.2 occurs, for any $u\in[0,\beta/4]$, if
$\sup_{s}\left|F_{S\mid\check{\theta}_{n}}\left(s\right)-\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}\left(s\right)\right|\leq 2f_{\min}u$
then $|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\leq u$.
Proof.
For simplicity, in the proof we denote $q_{p}\left(S\mid\check{\theta}_{n}\right)$ by $q_{p}$. By Proposition C.3, for $u\in[0,\beta/4]$, $|q_{\left(1-\alpha\right)_{m}}-\zeta-u|\leq 3\beta/4$ and $|q_{\left(1-\alpha\right)_{m}}-\zeta+u|\leq 3\beta/4$, i.e., $q_{\left(1-\alpha\right)_{m}}-u\in\mathcal{I}$ and $q_{\left(1-\alpha\right)_{m}}+u\in\mathcal{I}$ for $\mathcal{I}$ defined in Proposition C.2. Hence, in this case,
$F_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)\leq F_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}\right)-2f_{\min}u=\left(1-\alpha\right)_{m}-2f_{\min}u$
$F_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)\geq F_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}\right)+2f_{\min}u=\left(1-\alpha\right)_{m}+2f_{\min}u$
By assumption,
$\left|F_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)-\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)\right|\leq 2f_{\min}u$
$\left|F_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)-\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)\right|\leq 2f_{\min}u$
Then
$\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}-u\right)\leq\left(1-\alpha\right)_{m},\qquad\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}\left(q_{\left(1-\alpha\right)_{m}}+u\right)\geq\left(1-\alpha\right)_{m}$
Since $\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}$ is non-decreasing, we have
$\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right):=\inf\{u^{\prime}\in\mathcal{S}_{m}:\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}\left(u^{\prime}\right)\geq\left(1-\alpha\right)_{m}\}\in\left[q_{\left(1-\alpha\right)_{m}}-u,q_{\left(1-\alpha\right)_{m}}+u\right]$
where $\mathcal{S}_{m}$ is the set of scores of the calibration data. Then,
$|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\leq u.$
By the Dvoretzky–Kiefer–Wolfowitz Inequality (Lemma B.13),
$\mathbb{P}\left[\sup_{s}\left|F_{S\mid\check{\theta}_{n}}\left(s\right)-\hat{F}^{\left(m\right)}_{S\mid\check{\theta}_{n}}\left(s\right)\right|\geq 2f_{\min}u\right]\leq 2\exp{\left(-8mf_{\min}^{2}u^{2}\right)}$
Thus, by Lemma B.12, given that the event $V$ occurs,
$\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\geq u\;\Big|\;V\right]\leq 2\exp{\left(-8mf_{\min}^{2}u^{2}\right)},\quad u\in[0,\beta/4].$
Specifically,
$\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\geq\beta/4\;\Big|\;V\right]\leq 2\exp{\left(-8mf_{\min}^{2}(\beta/4)^{2}\right)}$
Then, for any $u>\beta/4$,
$\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\geq u\;\Big|\;V\right]\leq 2\exp{\left(-8mf_{\min}^{2}(\beta/4)^{2}\right)}$
Since $|S|\leq R$, $|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\leq 2R$. By the layer cake representation of the expectation of a non-negative random variable $Z$, which is $\mathbb{E}[Z]=\int_{0}^{\infty}\mathbb{P}[Z\geq u]\ du$,
$\mathbb{E}_{\check{\theta}_{n}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\;\Big|\;V\right]$
$=\int_{0}^{2R}\mathbb{P}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\geq u\;\Big|\;V\right]du$
$\leq\int_{0}^{\beta/4}2\exp{\left(-8mf_{\min}^{2}u^{2}\right)}du+\int_{\beta/4}^{2R}2\exp{\left(-8mf_{\min}^{2}(\beta/4)^{2}\right)}du$
$\leq 2\int_{0}^{\infty}\exp{\left(-8mf_{\min}^{2}u^{2}\right)}du+4R\exp{\left(-8f_{\min}^{2}(\beta/4)^{2}m\right)}$
$=\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2m}}+4R\exp{\left(-\frac{1}{2}f_{\min}^{2}\beta^{2}m\right)}$
Therefore, we have
$\mathbb{E}_{\check{\theta}_{n},\mathcal{D}_{\mathrm{cal}}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}_{n}\right)|\right]$
$\leq\mathbb{P}\left[V\right]\cdot\mathbb{E}_{\check{\theta}_{n}}\left[|\hat{q}_{\left(1-\alpha\right)_{m}}\left(S_{m}\mid\check{\theta}_{n}\right)-q_{\left(1-\alpha\right)_{m}}\left(S\mid\check{\theta}{n}\right)|\;\Big|\;V\right]+\mathbb{P}\left[V^{c}\right]\cdot 2R$
$\leq\frac{\sqrt{\pi}}{2f_{\min}\sqrt{2m}}+4R\exp{\left(-\frac{f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}}{8f_{\max}^{2}}m\right)}+2R\delta$
Picking $\delta=\frac{257\lambda_{\max}^{2}f_{\max}^{3}B^{2}d}{\lambda_{\min}^{4}f_{\min}^{2}\min\{\alpha^{2},(1-\alpha)^{2}\}n}$ completes the proof of Proposition C.4. ∎
Appendix D Additional Experiments on Synthetic Data
D.1 Data Generation in Section 6
Figure 4: The probability density function of $Y|X=x$ for synthetic dataset.
The sampler of the data distribution $\mathcal{P}$ is constructed as follows. A vector $\theta_{0}$ is first drawn from $\theta_{0}\sim\text{Uniform}([1,2]^{2})$. The covariate $X$ is sampled uniformly from $\mathcal{X}=[1,20]^{2}$, i.e., $X\sim\text{Uniform}([1,20]^{2})$. Then, the probability density function of the conditional distribution $Y|X=x$ is constructed over support $y_{\min},y_{\max}]$, where $y_{\text{max}}=[20,20]^{\top}\theta_{0}$ and $y_{\text{min}}=-y_{\text{max}}$. The conditional p.d.f., illustrated in Figure 4, is piecewise affine with five segments, symmetric about zero. The central segment carries probability mass $(1-\alpha_{0})$, and each the other four segments carries $\alpha_{0}/4$, where $\alpha_{0}=0.005$ is chosen to be smaller than the smallest miscoverage level considered in the experiments. The model is well-specified (Assumption 3.1) for $\gamma\in\{\alpha/2,1-\alpha/2\}$ and all $\alpha\in(\alpha_{0},{1}/{2})$ by taking $\theta^{*}(\gamma)=\frac{1-2(1-\gamma)}{1-\alpha_{0}}\theta_{0}$, and hence the true quantile functions $t_{\gamma}(x;\theta^{*}(\gamma))=\frac{1-2(1-\gamma)}{1-\alpha_{0}}\theta_{0}^{\top}x$. Then we can draw $y\sim Y|X=x$ from reject sampling to obtain $(x,y)$.
D.2 Validating Regime of $\mathcal{O}(1/(n\alpha^{2}))$
In the regime where $\alpha=o(n^{-1/4})$ and $\alpha=\omega(n^{-1/2})$, theory predicts that the length deviation should scale as $\mathcal{O}(1/(n\alpha^{2}))$, corresponding to the middle regime (green) in Figure 2. To validate this dependence, we pick $\alpha$ at several small values $\alpha=\{0.01,0.02,0.025,0.03\}$ and vary the training size $n$, plotting the length deviation against $1/(n\alpha^{2})$ on a log–log scale. The fitted regression line (red) in Figure 5 yields a slope of approximately $0.92$, which is close to the theoretical value of $1$. The empirical results support the predicted theoretical scaling, indicating the upper bound accurately captures the observed dependence.
Figure 5: Log–log regression of length deviation $\Delta$ versus $1/(n\alpha^{2})$ for relatively small $\alpha$.
Appendix E Experiments on Real-World Data
The Medical Expenditure Panel Survey (MEPS) Panels 19111https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-181 and 20222https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-192 are standard datasets used for benchmarking and comparative analysis in the quantile regression literature. These panels comprise 15,785, 17,541, and 15,656 samples, respectively. Each sample consists of 139 features, including 2 categorical features, 4 continuous features, and 133 boolean features. Throughout experiments, we train ridge regression models with stochastic gradient descent (SGD) optimizer with a step size tuned using successive halving for 1 epoch.
Conformalized Median Regression (CQR).
We examine the effect of the training set size $n$ and the calibration set size $m$ on the prediction set length using the MEPS’20 dataset, comparing the empirical results with the theoretical bound in Theorem 3.2. Since the oracle quantile interval length $|\mathcal{C}^{*}(X)|=q_{1-\alpha/2}(Y|X)-q_{\alpha/2}(Y|X)$ depends on $\alpha$, we evaluate the expected absolute deviation $\mathbb{E}[||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)||]$ for $\alpha\in[0.01,0.05,0.1,0.2]$. We reserve 20% of the dataset for testing length deviation. The remaining 80% was partitioned for training and calibration: the training size $n$ varied from 10% to 80% in increments of 10%, while the calibration $m$ was chosen from $5\%,10\%,15\%,20\%$} of the remaining data after allocating the training set. The results, shown in Fig. 6, confirm two key insights from our theoretical analysis. First, increasing the calibration set size $m$ reduces the expected length deviation. Second, for a fixed sample size, a larger miscoverage level $\alpha$ leads to a smaller deviation with lower variance, which aligns with the $\alpha$-dependence in the theoretical rate.
Conformalized Median Regression (CMR).
Figure 7 presents experimental results on length deviation for the MEPS’19 dataset under the CMR framework. The experimental setup mirrors that of the previous CQR analysis, adapted here for median regression. Consistent with Theorem 4.1, we observe that smaller values of $\alpha$ yield significantly larger length deviations.
Figure 6: The efficiency of CQR. Here the y-axis represents $\left||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)|\right|$, where the interval length $|\mathcal{C}^{*}(X)|$ is approximated by its estimate with same $\alpha$ and largest training and calibration sample sizes.
Figure 7: The efficiency of CMR. Here the y-axis represents $\left||\mathcal{C}(X)|-|\mathcal{C}^{*}(X)|\right|$, where the interval length $|\mathcal{C}^{*}(X)|$ is approximated by its estimate with same $\alpha$ and largest training and calibration sample sizes.